### v5

Abstract
In this paper we study the Spanning Tree Congestion problem, where we are given an undirected graph $G=(V,E)$ and are asked to find a spanning tree $T$ of minimum maximum congestion. Here, the congestion of an edge $e\in T$ is the number of edges $uv\in E$ such that the (unique) path from $u$ to $v$ in $T$ traverses $e$. We consider this well-studied NP-hard problem from the point of view of (structural) parameterized complexity and obtain the following results:
  • We resolve a natural open problem by showing that Spanning Tree Congestion is not FPT parameterized by treewidth (under standard assumptions). More strongly, we present a generic reduction which applies to (almost) any parameter of the form “vertex-deletion distance to class $\mathcal{C}$”, thus obtaining W[1]-hardness for more restricted parameters, including tree-depth plus feedback vertex set, or incomparable to treewidth, such as twin cover. Via a slight tweak of the same reduction we also show that the problem is NP-complete on graphs of modular-width $4$.
  • Even though it is known that Spanning Tree Congestion remains NP-hard on instances with only one vertex of unbounded degree, it is currently open whether the problem remains hard on bounded-degree graphs. We resolve this question by showing NP-hardness on graphs of maximum degree $8$.
  • Complementing the problem’s W[1]-hardness for treewidth, we formulate an algorithm that runs in time roughly ${(k+w)}^{\mathcal{O}(w)}$, where $k$ is the desired congestion and $w$ the treewidth, improving a previous argument for parameter $k+w$ that was based on Courcelle’s theorem. This explicit algorithm pays off in two ways: it allows us to obtain an FPT approximation scheme for parameter treewidth, that is, a $(1+\varepsilon)$-approximation running in time roughly ${(w/\varepsilon)}^{\mathcal{O}(w)}$; and it leads to an exact FPT algorithm for parameter clique-width+k+k via a Win/Win argument.
  • Finally, motivated by the problem’s hardness for most standard structural parameters, we present FPT algorithms for several more restricted cases, namely, for the parameters vertex-deletion distance to clique; vertex integrity; and feedback edge set, in the latter case also achieving a single-exponential running time dependence on the parameter.
keywords:
Parameterized Complexity, Treewidth, Graph Width Parameters
keywords:
linktext=https://arxiv.org/abs/2410.08314
category:
[…]
1 Introduction
One of the most well-studied types of problems in network optimization involves finding, for a given graph $G$, a spanning tree of $G$ that optimizes a certain objective. In this paper we focus on a well-known problem of this type called Spanning Tree Congestion. The motivation of this problem can be summarized as follows: Every edge $e$ of a spanning tree $T$ is selected with the goal of maintaining connectivity between the two parts of the graph given by the two components of TeT-e. We can then think of every other edge $e^{\prime}$ with endpoints in both components of TeT-e as being “simulated” by a path in $T$ that traverses $e$; hence, the more such edges exist, the more $e$ is used and “congested”. Our optimization goal, then, is to find a tree where all edges have congestion as low as possible, because in such a tree each selected edge is responsible for simulating only a small number of non-selected edges and therefore the tree can be thought of as a sparse approximate representation of the original graph. Equivalently, for a spanning tree $T$ of $G$, we say that the detour of an edge $\{u,v\}\in E(G)$ in $T$ is the unique uuvv path in $T$. The number of detours that traverse an edge in $T$ constitutes its congestion, while the congestion of $T$ is defined as the maximum over the congestion of all of its edges. [^2] This has also been referred to as the edge remember number of $G$ relative to $T$ in the literature [tcs/Bodlaender98, Section 11]. The spanning tree congestion of $G$, denoted by $\mathrm{stc}(G)$, is the minimum congestion over all of its spanning trees, and Spanning Tree Congestion asks, given $G$ and an integer $k$, whether $\mathrm{stc}(G)\leq k$.
Spanning trees of low congestion are a natural notion that is well-studied both from the combinatorial and the algorithmic point of view. Unsurprisingly, Spanning Tree Congestion is NP-complete [phd_thesis, Section 5.6]. It therefore makes sense to study the parameterized complexity of this problem, as parameterized complexity is one of the main tools for dealing with computational intractability. [^3] Throughout the paper we assume that the reader is familiar with the basics of parameterized complexity, as given in standard textbooks [books/CyganFKLMPPS15]. The most natural parameter one could consider is perhaps the objective value $k$, but unfortunately the problem is known to be NP-hard for all fixed $k\geq 5$ [algorithmica/BodlaenderFGOL12, algorithmica/LuuC24]. This motivates us to focus on structural graph parameters, where much less is currently known. Indeed, it is so far open whether Spanning Tree Congestion is fixed-parameter tractable for treewidth, which is the most widely studied parameter of this type (this is mentioned as an open problem in [birthday/Otachi20]). What is known, however, is that the problem is FPT when parameterized by both treewidth and $k$ [algorithmica/BodlaenderFGOL12] and that the problem is NP-hard on graphs of clique-width at most $3$ (implied by the NP-hardness on chain graphs [jgaa/OkamotoOUU11]).
$\mathrm{vc}$$\mathrm{vi}$$\mathrm{td}$$\mathrm{tw}$$\mathrm{cw}$$\mathrm{fes}$$\mathrm{fvs}$$\mathrm{nd}$$\mathrm{mw}$$\mathrm{tc}$$\mathrm{sb}$$\mathrm{dtc}$$\mathrm{cvd}$
Figure 1: Our results and hierarchy of the related graph parameters (see LABEL:paragraph:graph_parameters for their definitions). For any graph, if the parameter at the tail of an arrow is a constant, that is also the case for the one at its head. Green indicates that the problem is FPT (LABEL:thm:distance_to_clique, LABEL:thm:vi and 5.43), orange W[1]-hardness (Theorem˜1), and red para-NP-hardness (LABEL:thm:hardness:modularwidth). We additionally show fixed-parameter tractability by $\mathrm{cw}+k$ (LABEL:thm:cw), para-NP-hardness by maximum degree (Theorem˜3.19), as well as develop an FPT-AS for $\mathrm{tw}$ (LABEL:thm:tw-apx). Prior to this work, it was only known that the problem is FPT by $\mathrm{tw}+k$ [algorithmica/BodlaenderFGOL12] and para-NP-hard by clique-width [jgaa/OkamotoOUU11].
Our Contribution.
Our aim in this paper is to present a clarified and much more detailed picture of how the complexity of Spanning Tree Congestion depends on treewidth and other notions of graph structure (see Figure˜1 for a synopsis of our results).
We begin our work by considering the natural open problem we mentioned above, namely whether Spanning Tree Congestion is FPT parameterized by treewidth. We answer this question in the negative and indeed prove something much stronger: Let $\mathcal{C}$ be any class of graphs that satisfies the (very mild) requirement that for each integer $i$ there exists a connected graph in $\mathcal{C}$ that has $i$ vertices. Then, for any such class $\mathcal{C}$, Spanning Tree Congestion is W[1]-hard parameterized by the vertex deletion distance to a disjoint union of graphs belonging to $\mathcal{C}$. As a corollary, if we set $\mathcal{C}$ to be the class of all stars, Spanning Tree Congestion is shown to be W[1]-hard for parameter vertex-deletion distance to star-forest, hence also for parameter tree-depth plus feedback vertex set (and consequently also for treewidth). Alternatively, by setting $\mathcal{C}$ to be the class of all cliques, our proof establishes W[1]-hardness parameterized by the cluster vertex deletion number, and more strongly by the twin-cover of the input graph [dmtcs/Ganian15]. With a couple of modifications, we then show in LABEL:thm:hardness:modularwidth that Spanning Tree Congestion remains NP-complete even on graphs of modular-width at most $4$, linear clique-width $3$, and shrub-depth $2$, improving over the previously mentioned hardness result of Okamoto, Otachi, Uehara, and Uno for clique-width $3$ [jgaa/OkamotoOUU11]. As a matter of fact, the graphs resulting from our reduction are interval graphs, thus settling the open question posed by Otachi [birthday/Otachi20] on the polynomial-time solvability of Spanning Tree Congestion on interval graphs in the negative.
Moving on, we consider the tractability of the problem in graphs of constant degree. All previous NP-hardness results [algorithmica/BodlaenderFGOL12, algorithmica/LuuC24] require at least one vertex of unbounded degree. However, assuming that the graph has bounded degree seems potentially algorithmically useful, as recent work by Kolman [iwoca/Kolman24] shows that instances of polylogarithmic maximum degree are amenable to a polynomial approximation algorithm of ratio $o(n)$ (this is non-trivial, as the best known ratio on general graphs is $n/2$, trivially achieved by any spanning tree [birthday/Otachi20]). [^4] A very recent work by Kolman [stacs/Kolman25] improves over this and presents a polynomial-time approximation algorithm of ratio ${\cal O}(\Delta\cdot\log^{3/2}n)$, where $\Delta$ denotes the maximum degree of the graph. Our next result is to answer an open question posed by Kolman [iwoca/Kolman24] and show that the problem in fact remains NP-hard even on graphs of degree at most $8$ (Theorem˜3.19). To this end, we make use of a novel gadget based on grids, simulating the double-weighted edges introduced by Luu and Chrobak [algorithmica/LuuC24].
Coming back to treewidth, we recall that Bodlaender, Fomin, Golovach, Otachi, and van Leeuwen [algorithmica/BodlaenderFGOL12] showed that, when $k$ is part of the parameter, Spanning Tree Congestion is expressible in MSO2 logic, thus due to Courcelle’s theorem [iandc/Courcelle90] fixed-parameter tractable by $\mathrm{tw}+k$. We improve upon this by providing an explicit FPT algorithm of running time ${(\mathrm{tw}+k)}^{{\cal O}(\mathrm{tw})}n^{{\cal O}(1)}$. In addition to providing a concrete reasonable upper-bound on the running time (which cannot be done with Courcelle’s theorem), this explicit algorithm allows us to obtain two further interesting extensions. First, using a technique introduced by Lampis [icalp/Lampis14], we develop an efficient FPT approximation scheme (FPT-AS) when parameterized solely by $\mathrm{tw}$, that is, a $(1+\varepsilon)$-approximate algorithm running in time ${(\mathrm{tw}/\varepsilon)}^{{\cal O}(\mathrm{tw})}n^{{\cal O}(1)}$; notice that an efficient FPT-AS is the best we can hope for in this setting, given the W[1]-hardness following from Theorem˜1. Second, using a Win/Win argument based on a result of Gurski and Wanke [wg/GurskiW00], we lift our algorithm to also show an explicit FPT algorithm for the more general parameter $\mathrm{cw}+k$, where $\mathrm{cw}$ denotes the clique-width of the input graph.
Finally, given all the previously mentioned hardness results, we next aim to determine which structural parameters do render the problem fixed-parameter tractable. As a consequence of Theorem˜1, the problem remains intractable even on very restricted (dense and sparse) graph classes, we must therefore focus on parameters that evade this hardness result. We consider three cases: First, the parameter “distance to clique” is not covered by Theorem˜1 because the graph obtained after removing the deletion set has one component; we show in LABEL:thm:distance_to_clique that Spanning Tree Congestion is FPT in this case. Second, the parameter vertex integrity is not covered by Theorem˜1, as all components of the graph obtained after removing the deletion set have bounded size; we show in LABEL:thm:vi that Spanning Tree Congestion is FPT in this case as well, via a reduction to an ILP with an FPT number of variables. Third, we consider the parameter feedback edge set, which also falls outside the scope of Theorem˜1, and obtain a linear kernel, which leads to an FPT algorithm with single-exponential parameter dependence for this case.
Related Work.
Spanning Tree Congestion was formally introduced by Ostrovskii [dm/Ostrovskii04], though it had also been previously studied under a different name [mst/Simonson87]. There is a plethora of graph-theoretical results in the literature [dmgt/CastejonO09, dm/Hruska08, dmgt/KozawaO11, dm/KozawaOY09, dm/Law09, dm/LowensteinRR09, dm/Ostrovskii10], as well as some algorithmic ones [algorithmica/BodlaenderFGOL12, dm/BodlaenderKMO11, icalp/ChandranCI18, jgaa/OkamotoOUU11]. See also the survey of Otachi [birthday/Otachi20]. Spanning Tree Congestion is known to be polynomial-time solvable if $k\leq 3$ [algorithmica/BodlaenderFGOL12], and NP-hard for all fixed $k\geq 5$ [algorithmica/LuuC24]; the case $k=4$ remains open. Okamoto et al. [jgaa/OkamotoOUU11] have presented an algorithm running in time $2^{n}n^{{\cal O}(1)}$, improving over the brute-force one. Regarding specific graph classes, it is known to be polynomial-time solvable for outerplanar graphs [dm/BodlaenderKMO11], two-dimensional Hamming graphs [dmgt/KozawaO11], complete $k$-partite graphs, and two-dimensional tori [dm/KozawaOY09]. On the other hand, it is NP-hard for planar, split, and chain graphs [algorithmica/BodlaenderFGOL12, jgaa/OkamotoOUU11], with the latter result implying NP-hardness for graphs of clique-width at most $3$. For fixed $k$, the problem is expressible in MSO2 logic [algorithmica/BodlaenderFGOL12], thus due to standard metatheorems [iandc/Courcelle90] it is FPT by $\mathrm{tw}+k$. Kozawa, Otachi, and Yamazaki [dm/KozawaOY09] showed a combinatorial bound (then improved in [algorithmica/BodlaenderFGOL12]) which proves that for all graphs $G$, $\mathrm{tw}(G)={\cal O}(\mathrm{stc}(G)\Delta(G))$, where $\Delta$ denotes the maximum degree of the input graph. Combining these, Bodlaender et al. [algorithmica/BodlaenderFGOL12] show that Spanning Tree Congestion is FPT by $\Delta+k$, as well as that it is solvable in polynomial time for fixed $k$ on apex-minor-free graphs. There are also some results regarding the problem’s approximability [algorithmica/BodlaenderFGOL12, iwoca/Kolman24, stacs/Kolman25, algorithmica/LuuC24].
Finding a spanning tree $T$ of a connected graph such that $T$ adheres to some constraint, i.e., $T\in\mathcal{T}$ for some family of trees $\mathcal{T}$, is an interesting combinatorial question in its own right, that oftentimes finds applications to other algorithmic problems. Examples of studied properties include trees of maximum number of branch or leaf vertices [siamdm/BonsmaZ11, siamdm/DeBiasioL19, mst/FellowsLMMRS09, icalp/GarganoHSV02, mfcs/GarganoR23, siamdm/KleitmanW91, algorithmica/Lampis12], of minimum maximum degree [arxiv/BojikianFGHS25, jacm/SinghL15], and others [siamcomp/Agarwal92, dam/BercziKKYY25, jcss/BergougnouxBFGRS25, algorithmica/HalldorssonKMT21, NageleZ19]. One such important variant of Spanning Tree Congestion is the Tree Spanner problem [siamdm/CaiC95], where one asks for a spanning tree of minimum stretch. The latter has been extensively studied [siamcomp/AbrahamN19, swat/BorradaileCEMN20, jcss/DraganFG11, siamcomp/ElkinEST08, siamcomp/EmekP08, dam/FeketeK01, ipl/FominGL11], and the two problems are known to be tightly connected, especially on planar graphs [algorithmica/LuuC24, birthday/Otachi20].
Lastly, a closely-related structural graph parameter is the so-called edge-cut width [wg/BrandCGHK22] or local feedback edge number [nips/GanianK21]. This is, roughly speaking, the vertex variant of spanning tree congestion, where one asks to minimize the maximum congestion over the vertices of the spanning tree, where the congestion of a vertex is defined as the number of detours containing it. [^5] Analogously, this has been referred to as the vertex remember number of $G$ relative to $T$ in the literature [tcs/Bodlaender98, Section 11]. Those parameters have been recently used in the setting of parameterized complexity to show various tractability and incompressibility results [wg/BrandCGHK22, nips/GanianK21], and we believe that our work might provide insights into the parameterized (in)tractability of their computation.
Organization.
In Section˜2 we discuss the general preliminaries. Subsequently, in Section˜3 we present the various hardness results, followed by Section˜4 where we present the explicit FPT algorithm when parameterized by $\mathrm{tw}+k$, as well as the two results that make use of this. Moving on, in Section˜5 we present various fixed-parameter tractability results. Lastly, in Section˜6 we present the conclusion as well as some directions for future research.
2 Preliminaries
Throughout the paper we use standard graph notation [Diestel17], and we assume familiarity with the basic notions of parameterized complexity [books/CyganFKLMPPS15]. All graphs considered are undirected without loops. For a graph $G=(V,E)$ and $S\subseteq V$, we denote the open neighborhood of $S$ by $N_{G}(S)=(\bigcup_{s\in S}N_{G}(s))\setminus S$. For $x,y\in\mathbb{Z}$, let $x,y]=\{z\in\mathbb{Z}\mid x\leq z\leq y\}$, while $[x]=[1,x]$.
Let $G=(V,E)$ be a connected graph and $T$ a spanning tree of $G$. The detour for $\{u,v\}\in E$ in $T$ is the unique uuvv path in $T$. Note that the detour of $e\in E(T)$ is $e$ itself. The congestion of $e\in E(T)$, denoted $\mathrm{cng}_{G,T}(e)$, is the number of edges in $G$ whose detours contain $e$. In other words, $\mathrm{cng}_{G,T}(e)$ is the size of the fundamental cutset of $T$ defined by $e$, that is, $\mathrm{cng}_{G,T}(e)=|E(V(T_{e,1}),V(T_{e,2}))|$, where $E(X,Y)=\{\{x,y\}\in E\mid x\in X,\,y\in Y\}$ for disjoint vertex sets $X,Y\subseteq V$ and $T_{e,1}$ and $T_{e,2}$ are the two subtrees of $T$ obtained by cutting $e$. The congestion of $T$, denoted $\mathrm{cng}_{G}(T)$, is defined as the maximum over the congestion of all edges in $T$, i.e., $\mathrm{cng}_{G}(T)=\max_{e\in E(T)}\mathrm{cng}_{G,T}(e)$. The spanning tree congestion of $G$, denoted $\mathrm{stc}(G)$, is the minimum congestion over all spanning trees of $G$. Given a connected graph $G$ and an integer $k\in\mathbb{Z}^{+}$, Spanning Tree Congestion asks to determine whether $\mathrm{stc}(G)\leq k$.
Let $G=(V,E)$ be a graph. The vertex integrity of $G$, denoted by $\mathrm{vi}(G)$, is the minimum integer $k$ such that there is a vertex set $S\subseteq V$ with $|S|+\max_{C\in\mathtt{cc}(G-S)}|V(C)|\leq k$, where $\mathtt{cc}(G-S)$ denotes the set of connected components in GSG-S. The twin-cover number of $G$, denoted by $\mathrm{tc}(G)$, is the size of the smallest vertex set (called twin-cover) whose removal results in a cluster graph, with the constraint that each clique is composed of true twins in $G$ [dmtcs/Ganian15]. If we drop the constraint, this is the cluster vertex deletion number [mfcs/DouchaK12], denoted by $\mathrm{cvd}(G)$. The modular-width of $G$ ([GajarskyLMMO15, GajarskyLO13]) is the smallest integer $k$ such that, either $|V|\leq k$, or $V$ can be partitioned into at most $k^{\prime}\leq k$ sets $V_{1},\ldots,V_{k^{\prime}}$, with the following two properties: (i) for all $i\in[k^{\prime}]$, $V_{i}$ is a module of $G$, (ii) for all $i\in[k^{\prime}]$, $G[V_{i}]$ has modular width at most $k$. For the definition of the rest of the parameters appearing in Figure˜1 as well as known relations between them, we refer to LABEL:paragraph:graph_parameters.
3 Hardness results
In this section we present various hardness results for Spanning Tree Congestion. We start with showing in Section˜3.1 that the problem is W[1]-hard parameterized by the distance to the disjoint union of graphs in $\mathcal{C}$, for any family of graphs $\mathcal{C}$ that contains connected graphs of any order. Moving on, in Section˜3.2 we adapt our proof and prove NP-hardness for graphs of modular-width at most $4$. Finally, in Section˜3.3 we introduce a novel gadget simulating the double-weighted edges previously used by Luu and Chrobak [algorithmica/LuuC24], of which we make use of in order to show NP-hardness for graphs of constant maximum degree.
3.1 Distance to disjoint union
We start by stating the main theorem of this subsection.
Theorem 1.
Spanning Tree Congestion is W[1]-hard parameterized by vertex-deletion distance to disjoint union of graphs in $\mathcal{C}$, where $\mathcal{C}$ is any graph class such that, for all $p\in\mathbb{Z}^{+}$, $\mathcal{C}$ contains a connected $p$-vertex graph which can be generated in time $p^{{\cal O}(1)}$.
By taking the set of stars as $\mathcal{C}$, Theorem˜1 implies the W[1]-hardness parameterized by distance to star forest.
Corollary 3.2.
Spanning Tree Congestion is W[1]-hard parameterized by distance to star forest (and thus by tree-depth ++ feedback vertex set number).
If $\mathcal{C}$ is the set of complete graphs, Theorem˜1 implies W[1]-hardness parameterized by cluster vertex deletion number [mfcs/DouchaK12]. In fact, as we will later see, the proof of Theorem˜1 more strongly implies W[1]-hardness parameterized by twin-cover number [dmtcs/Ganian15].
Corollary 3.3.
Spanning Tree Congestion is W[1]-hard parameterized by twin-cover number.
For an edge-weighted graph $G=(V,E;\mathrm{w})$ with $\mathrm{w}\colon E\to\mathbb{Z}^{+}$, we define its spanning tree congestion by setting the congestion of an edge $e$ in a spanning tree $T$ of $G$ as $\mathrm{cng}_{G,T}(e)=\mathrm{w}(E(V(T_{e,1}),V(T_{e,2})))$, where $\mathrm{w}(F)=\sum_{e\in F}\mathrm{w}(e)$ for $F\subseteq E$. The following proposition provides a connection between the weighted and the unweighted case.
Proposition 3.4 ([algorithmica/BodlaenderFGOL12]).
Let $G=(V,E;\mathrm{w})$ be an edge-weighted graph and $G^{\prime}=(V^{\prime},E^{\prime})$ be an unweighted graph obtained from $G$ by replacing each weighted edge $\{u,v\}$ with $\mathrm{w}(\{u,v\})$ internally vertex-disjoint uuvv paths of any lengths, where one of them may be $\{u,v\}$ itself. Then, $\mathrm{stc}(G)=\mathrm{stc}(G^{\prime})$.
In the following, we prove Theorem˜1 by a reduction from Unary Bin Packing. Given unary encoded integers $t,a_{1},\dots,a_{n}\in\mathbb{Z}^{+}$ with $\sum_{i\in[n]}a_{i}=tB$, Unary Bin Packing asks whether there is a partition $(A_{1},\dots,A_{t})$ of $[n]$ such that $\sum_{i\in A_{j}}a_{i}=B$ for each $j\in[t]$. It is known that Unary Bin Packing is W[1]-hard parameterized by $t$ [jcss/JansenKMS13]. We assume that $t\geq 3$ since otherwise the problem can be solved in polynomial time. We now proceed to presenting the reduction.
Construction.
Let $\mathcal{I}=\langle t;a_{1},\dots,a_{n}\rangle$ be an instance of Unary Bin Packing with $\sum_{i\in[n]}a_{i}=tB$. For each $i\in[n]$, let $G_{i}=(V_{i},E_{i})$ be a connected $a_{i}$-vertex graph belonging to $\mathcal{C}$. We set k=5(t1)Bk=5(t-1)B and construct an edge-weighted graph $G=(V,E;\mathrm{w})$ as follows.
  1. 1.
    Take the disjoint union of all $G_{i}$ for $i\in[n]$.
  2. 2.
    Add a set of vertices $Q=\{v_{j}\mid j\in[t]\}$ and all possible edges between $Q$ and $\bigcup_{i\in[n]}V_{i}$.
  3. 3.
    Add a vertex $r$ and all possible edges between $r$ and $Q$.
  4. 4.
    Set $\mathrm{w}(\{r,v_{j}\})=3(t-1)B$ for $j\in[t]$, and $\mathrm{w}(e)=1$ for all other edges.
Proposition˜3.4 implies that we can construct, in time polynomial in $n$ and $B$, an unweighted graph $G^{\prime}$ from $G$ such that $\mathrm{stc}(G)=\mathrm{stc}(G^{\prime})$, where each weighted edge $e$ of weight $\mathrm{w}(e)\geq 2$ in $G$ is replaced by $\mathrm{w}(e)$ internally vertex-disjoint paths of length $2$ between the endpoints of $e$. Observe that $G^{\prime}-(\{r\}\cup Q)$ is the disjoint union of all $G_{i}$ and the singleton components that correspond to the middle vertices of the paths replacing weighted edges. Since the single-vertex graph belongs to $\mathcal{C}$, $G^{\prime}$ has distance $|\{r\}\cup Q|=t+1$ to disjoint union of graphs in $\mathcal{C}$. We can also see that if $\mathcal{C}$ is the set of complete graphs, then $\{r\}\cup Q$ is a twin-cover. Thus, to prove Theorem˜1 (together with Corollary˜3.3), it suffices to show that $\mathrm{stc}(G)\leq k$ if and only if $\mathcal{I}$ is a yes-instance of Unary Bin Packing.
Lemma 3.5.
If $\mathcal{I}$ is a yes-instance of Unary Bin Packing, then $\mathrm{stc}(G)\leq k$.
Proof 3.6.
Let $(A_{1},\dots,A_{t})$ be a partition of $[n]$ such that $\sum_{i\in A_{j}}a_{i}=B$ for each $j\in[t]$. We construct a spanning tree $T$ of $G$ by setting
$E(T)=\{\{r,v_{j}\}\mid j\in[t]\}\cup\bigcup_{j\in[t]}\{\{u,v_{j}\}\mid u\in V_{i},\,i\in A_{j}\}.$
Each edge $\{u,v_{j}\}\in E(T)$ with $u\in V_{i}$ has congestion $\deg_{G}(u)=t+\deg_{G_{i}}(u)\leq t+a_{i}-1<k$ since $u$ is a leaf of $T$.
For each $v_{j}$, let $S_{j}$ be the set of vertices of the component of $T-\{r,v_{j}\}$ containing $v_{j}$. By the construction, we can see that $S_{j}=\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. Thus,
$\mathrm{cng}_{G,T}(\{r,v_{j}\})$ $=\mathrm{w}(E(S_{j},V\setminus S_{j}))$
$=\mathrm{w}(\{r,v_{j}\})+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[n]\setminus A_{j}}V_{i})|+|E(\textstyle\bigcup_{i\in A_{j}}V_{i},\;Q\setminus\{v_{j}\})|$
$=3(t-1)B+(t-1)B+(t-1)B=k.$
Lemma 3.7.
If $\mathrm{stc}(G)\leq k$, then $\mathcal{I}$ is a yes-instance of Unary Bin Packing.
Proof 3.8.
Let $T$ be a spanning tree of $G$ with congestion at most $k$.
[…]
Next we show that for each $i\in[n]$, there exists exactly one index $j\in[t]$ such that at least one vertex in $V_{i}$ is adjacent to $v_{j}$ in $T$.
Claim 1.
For all $i\in[n]$, there exists $j\in[t]$ such that $N_{T}(V_{i})\cap Q=\{v_{j}\}$.
{claimproof}
There is at least one such $j\in[t]$ since $T$ is a spanning tree, i.e., $N_{T}(V_{i})\cap Q\neq\varnothing$. Suppose to the contrary that for some $h\in[n]$, there are two or more vertices in $Q$ that have neighbors in $V_{h}$. Since $V_{h}$ induces a connected subgraph of $G$ (i.e., $G_{h}=(V_{h},E_{h})$) and each edge $\{r,v_{j}\}$ is included in $T$, there is at least one edge $e_{h}\in E_{h}$ such that the detour for $e_{h}$ in $T$ contains $r$. Let $R=\sum_{j\in[t]}\mathrm{cng}_{G,T}(\{r,v_{j}\})$. The edge $e_{h}$ contributes $2$ to $R$ as its detour passes through two edges incident to $r$. Each edge $\{r,v_{j}\}$ contributes $\mathrm{w}(\{r,v_{j}\})$ to $R$. Now, for $u\in\bigcup_{i\in[n]}V_{i}$, let $j_{u}\in[t]$ be the index such that $v_{j_{u}}$ appears in the uurr path $P_{u,r}$ in $T$. Since $\{r,v_{j}\}\in E(T)$ for each $j\in[t]$, such $j_{u}$ is unique and $v_{j_{u}}$ appears right before $r$ in $P_{u,r}$. Observe that for each $j\in[t]\setminus\{j_{u}\}$, the detour for $\{u,v_{j}\}\in E$ in $T$ consists of $P_{u,r}$ and $v_{j}$, where $v_{j}$ appears right after $r$. This detour contributes $1$ to the congestion of each of the edges $\{r,v_{j_{u}}\}$ and $\{r,v_{j}\}$. The discussion so far implies that $R>tk$ as follows:
$R$ $\geq 2+\sum_{j\in[t]}\mathrm{w}(\{r,v_{j}\})+\sum_{u\in\bigcup_{i\in[n]}V_{i}}2(t-1)=2+3(t-1)Bt+2(t-1)tB=2+tk.$
This contradicts the assumption that each edge of $T$ has congestion at most $k$ and thus $R\leq kt$.
For $j\in[t]$, let $A_{j}=\{i\mid\exists u\in V_{i},\,\{u,v_{j}\}\in E(T)\}$. Claim˜1 implies that $(A_{1},\dots,A_{t})$ is a partition of $[n]$. In particular, the set of vertices of the component of $T-\{r,v_{j}\}$ containing $v_{j}$ is $\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. This implies that
$\mathrm{cng}_{G,T}(\{r,v_{j}\})$ $=\mathrm{w}(\{r,v_{j}\})+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[n]\setminus A_{j}}V_{i})|+|E(\textstyle\bigcup_{i\in A_{j}}V_{i},\;Q\setminus\{v_{j}\})|$
$=3(t-1)B+\sum_{i\in[n]\setminus A_{j}}a_{i}+(t-1)\sum_{i\in A_{j}}a_{i}$
$=3(t-1)B+tB+(t-2)\sum_{i\in A_{j}}a_{i}.$
Combining this with the assumption $\mathrm{cng}_{G,T}(\{r,v_{j}\})\leq k=5(t-1)B$, we obtain that $\sum_{i\in A_{j}}a_{i}\leq B$. (Recall that $t\geq 3$.) Since $\sum_{i\in[n]}a_{i}=tB$, we have $\sum_{i\in A_{j}}a_{i}=B$ for each $j\in[t]$.
3.2 Modular-width
In this subsection we consider Spanning Tree Congestion parameterized by the modular-width of the input graph [GajarskyLO13]. We prove that the problem remains NP-complete even on graphs of modular-width at most $4$; there is no graph of modular-width $3$ since there is no prime graph with three vertices, therefore our result leaves open the case of graphs of modular-width at most $2$. Our result, along with the fact that graphs of modular-width at most $4$ have clique-width at most $3$ (LABEL:thm:mw_and_cw), improves upon previous work by Okamoto et al. [jgaa/OkamotoOUU11], who showed that the problem is NP-hard for graphs of clique-width at most $3$. As a matter of fact, the graph we construct has linear clique-width [siamdm/FellowsRRS09] at most $3$, thus directly improving over said result. Nevertheless, LABEL:thm:mw_and_cw is an interesting result in its own right which we were not able to find in the literature. Furthermore, the family of graphs constructed by our reduction has shrub-depth at most $2$ [lmcs/GanianHNOM19, mfcs/GanianHNOMR12], therefore yielding para-NP-hardness for this parameterization as well. Importantly, the graphs resulting from our reduction are interval graphs, thus answering a previously posed open question on the polynomial-time solvability of Spanning Tree Congestion on interval graphs in the negative [birthday/Otachi20].
Theorem 3.9.
thm:mw_and_cw Any graph of modular-width at most $4$ has clique-width at most $3$.
Proof 3.10.
Observe that a graph $G$ with at most four vertices has clique-width at most $3$: if $G\neq P_{4}$, then it is a cograph and thus $\mathrm{cw}(G)\leq 2$; if $G=P_{4}$, then $\mathrm{cw}(G)=3$ as it is a tree but not a cograph (see e.g., [dam/CourcelleO00]).
Let $G$ be a graph of modular-width at most $4$, consisting of modules $M_{1},\ldots,M_{p}$ with $p\leq 4$, where each module $M_{i}$ induces a graph of clique-width at most $3$. Let $H$ be the quotient graph of $G$; that is, $H$ is the graph obtained from $G$ by replacing each module with a single vertex. For $i\in[p]$, let $v_{i}\in V(H)$ be the vertex corresponding to $M_{i}$. Since $H$ has at most four vertices, it admits a clique-width expression with at most three labels. From the expression of $H$, we construct $G$ by replacing the introduction of $v_{i}$ with label $\ell\in[3]$ with the introduction of $M_{i}$ with label $\ell$. (By “introducing $M_{i}$ with label $\ell$”, we mean the following steps: (1) construct $M_{i}$ using at most three labels; (2) relabel all vertices in $M_{i}$ by the label $\ell$.)
In the following, we mostly follow the proof of Theorem˜1 by setting $\mathcal{C}$ to be the class of all complete graphs, albeit with a few adaptations. First, the starting point of our reduction is the strongly NP-complete $3$-Partition problem. Second, we notice that even though the edge-weighted graph $G$ produced in the proof of Theorem˜1 has modular-width $2$ when $\mathcal{C}$ is the class of complete graphs (let one module contain the vertices of $Q$ and the other the rest), that is not the case for the unweighted graph $G^{\prime}$ produced by Proposition˜3.4. In order to overcome this bottleneck, we emulate the weights of the edges by introducing a sufficiently large clique in our construction.
Given $3n$ unary encoded (not necessarily distinct) integers $a_{i}\in\mathbb{Z}^{+}$ for $i\in[3n]$, where $\sum_{i\in[3n]}a_{i}=nB$ and $B/4<a_{i}<B/2$ for all $i\in[3n]$, $3$-Partition asks whether there is a partition $(A_{1},\ldots,A_{n})$ of $[3n]$ such that $\sum_{i\in A_{j}}a_{i}=B$ for all $j\in[n]$; notice that the bounds on the values of $a_{i}$ imply that if $\sum_{i\in A_{j}}a_{i}=B$, then $|A_{j}|=3$. It is well-known that $3$-Partition is strongly NP-complete [books/fm/GareyJ79, Theorem 4.4].
Theorem 3.11.
thm:hardness:modularwidth Spanning Tree Congestion is NP-complete on interval graphs of modular-width at most $4$, linear clique-width at most $3$, and shrub-depth at most $2$.
Proof 3.12.
Let $\mathcal{I}=\langle a_{1},\dots,a_{3n}\rangle$ be an instance of $3$-Partition with $\sum_{i\in[3n]}a_{i}=nB$ and $B/4<a_{i}<B/2$ for all $i\in[3n]$. Assume without loss of generality that $n\geq 3$. For each $i\in[3n]$, let $G_{i}=(V_{i},E_{i})$ be a clique on $a_{i}$ vertices, i.e., $G_{i}$ is a $K_{a_{i}}$. We construct the graph $G$ as follows:
  1. 1.
    Take the disjoint union of all $G_{i}$ for $i\in[3n]$.
  2. 2.
    Add a set of vertices $Q=\{v_{j}\mid j\in[n]\}$ that form a clique and all possible edges between $Q$ and $\bigcup_{i\in[3n]}V_{i}$.
  3. 3.
    Add a set of vertices $C=\{c_{z}\mid z\in[M]\}$ that form a clique and all possible edges between $C$ and $Q$, where M=4(n1)B+2(n1)M=4(n-1)B+2(n-1).
  4. 4.
    Add a vertex $w$, and all possible edges between $w$ and $C$.
This concludes the construction of $G$; we refer to Figure˜2 for an overview of the construction. We set k=M+2(n1)B+(n1)=3M/2k=M+2(n-1)B+(n-1)=3M/2 and we claim that $\mathcal{I}$ is a yes-instance of $3$-Partition if and only if $\mathrm{stc}(G)\leq k$. Before proving this, we provide upper bounds on $G$’s modular-width and linear clique-width. Furthermore, we prove that the class of graphs produced by our reduction has shrub-depth at most $2$.
Figure 2: The graph $G$ along with an interval representation of it.
Lemma 3.13.
$G$ has modular-width at most $4$ and linear clique-width at most $3$.
Proof.
For the modular-width, consider a modular decomposition of $G$ with modules containing vertex sets $\{w\}$, $C$, $Q$, and $\bigcup_{i\in[3n]}V_{i}$ respectively. As for the linear clique-width, we introduce the vertices of $G$ as follows:
  • first introduce [^6] When we say that we introduce a clique using labels $i$ and $j$, we first introduce one of its vertices with label $i$, and then as long as there is an unintroduced vertex of the clique, we introduce it with label $j$, join iji-j, and relabel $j\to i$. $G_{1}$ using labels $1$ and $2$,
  • for $i\in[2,3n]$, (i) introduce $G_{i}$ using labels $2$ and $3$; (ii) relabel $2\to 1$,
  • introduce $C$ using labels $2$ and $3$,
  • introduce vertex $w$ with label $3$, and join 2233,
  • relabel $2\to 1$,
  • for each vertex $v\in Q$, (i) introduce $v$ using label $2$; (ii) join 1122; (iii) relabel $2\to 1$.∎
Lemma 3.14.
The class of graphs $\mathcal{G}$ produced by our reduction has shrub-depth at most $2$.
Proof.
We will present a tree-model $(T,S)$ of four colors and depth $2$ for any graph $G\in\mathcal{G}$. Fix a graph $G\in\mathcal{G}$, and let $T$ be the rooted tree whose root is connected to the centers of $n+1$ stars, such that
  • there is a star per graph $G_{i}$, whose leaves are $G_{i}$’s vertices, colored with color $1$,
  • there is an additional star, whose leaves are the rest of the vertices of the graph, namely the vertices belonging to $Q$ and $C$ as well as vertex $w$. Color the vertices belonging to $Q$ with color $2$, those belonging to $C$ with color $3$, and $w$ with color $4$.
Finally, let $S=\{(1,1,1),\,(1,2,2),\,(2,1,2),\,(2,2,1),\,(3,3,1),\,(2,3,1),\,(3,2,1),\,(3,4,1),\,(4,3,1)\}$. Since every graph belonging to $\mathcal{G}$ admits a tree-model of four colors and depth $2$, the statement follows. ∎
Lemma 3.15.
If $\mathcal{I}$ is a yes-instance of $3$-Partition, then $\mathrm{stc}(G)\leq k$.
Proof.
Let $(A_{1},\dots,A_{n})$ be a partition of $[3n]$ such that $|A_{j}|=3$ and $\sum_{i\in A_{j}}a_{i}=B$ for all $j\in[n]$. We construct a spanning tree $T$ of $G$ by setting
$E(T)=\{\{c_{1},u\}\mid u\in(Q\cup C\cup\{w\})\setminus\{c_{1}\}\}\cup\bigcup_{j\in[n]}\{\{u,v_{j}\}\mid u\in V_{i},\,i\in A_{j}\}.$
Each edge $\{u,v_{j}\}\in E(T)$ with $u\in V_{i}$ has congestion $\mathrm{cng}_{G,T}(\{u,v_{j}\})=\deg_{G}(u)=n+a_{i}-1<k$ since $u$ is a leaf of $T$. Furthermore, all vertices $u\in(C\cup\{w\})\setminus\{c_{1}\}$ are leaves in $T$, thus $\mathrm{cng}_{G,T}(\{c_{1},u\})=\deg_{G}(u)\leq M+n<k$.
For each $v_{j}$, let $S_{j}$ be the set of vertices of the component of $T-\{c_{1},v_{j}\}$ containing $v_{j}$. By the construction, we can see that $S_{j}=\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. Thus,
$\mathrm{cng}_{G,T}(\{c_{1},v_{j}\})$ $=|E(S_{j},V(G)\setminus S_{j})|$
$=|E(\{v_{j}\},C)|+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[3n]\setminus A_{j}}V_{i})|+|E(\textstyle\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i},\;Q\setminus\{v_{j}\}\})|$
$=M+(n-1)B+(n-1)+(n-1)B=k.\qed$
Lemma 3.16.
If $\mathrm{stc}(G)\leq k$, then $\mathcal{I}$ is a yes-instance of $3$-Partition.
Proof.
Let $T$ be a spanning tree of $G$ with $\mathrm{cng}_{G}(T)\leq k$. We first prove that $T-(\bigcup_{i\in[3n]}V_{i})$ is a star centered on a vertex of $C$.
Claim 2.
There exists $\ell\in[M]$ such that $\{c_{\ell},u\}\in E(T)$, for all $u\in(Q\cup C\cup\{w\})\setminus\{c_{\ell}\}$.
{claimproof}
Let $Q\cup C\cup\{w\}$ denote the set of important vertices of $G$. Since $T$ is a tree, there exists a vertex $u\in V(G)$ such that each connected component of TuT-u has size at most $|V(G)|/2$. We will first prove that every connected component of TuT-u contains at most one important vertex.
Assume otherwise, and let $T_{1}$ denote a component of TuT-u containing more than one important vertices, with $e$ being the single edge connecting $u$ and $T_{1}$. Let $T_{2}$ denote the component of TeT-e containing $u$. Notice that $T_{1}$ and $T_{2}$ define a partition of the vertices of $C$, where $C_{i}=C\cap V(T_{i})$ for $i\in[2]$.
We show that $2\leq|C_{1}|<|C|-2=M-2$. If $C_{1}=\varnothing$, then $\mathrm{cng}_{G,T}(e)\geq 2M>k$, due to the edges between the (at least $2$) important vertices of $T_{1}$ and all vertices of $C_{2}=C$. If $|C_{1}|=1$, then $\mathrm{cng}_{G,T}(e)\geq 2(M-1)>k$ in an analogous way. Lastly, $|C_{1}|\leq|V(T_{1})|\leq|V(G)|/2=(nB+n+M+1)/2<M-2$.
In that case, it holds that
$\mathrm{cng}_{G,T}(e)$ $\geq|E(C_{1},C_{2})|=|C_{1}|\cdot(M-|C_{1}|)$
$\geq 2\cdot(M-2)$
$>\frac{3M}{2}=k,$
a contradiction.
Consequently, every connected component of TuT-u contains at most one important vertex. In that case, it follows that
  • either $u=w$, that is, every connected component of TwT-w contains at most one important vertex,
  • or $N_{T}(w)=\{u\}$, as every neighbor of $w$ is an important vertex.
Assume the first case, and notice that then, every vertex in the connected component of TwT-w that contains $v_{j}\in Q$ belongs to $Q\cup\bigcup_{i\in[3n]}V_{i}$. Since none of these vertices has an edge with $w$, $T$ is disconnected, a contradiction. The discussion so far implies that there exists $\ell\in[M]$ such that $c_{\ell}\in C\cap N_{T}(w)$ denotes the single neighbor of $w$ in $T$, with every connected component of $T-c_{\ell}$ having at most one important vertex.
To finish our proof, assume there exists $u\in(Q\cup C)\setminus\{c_{\ell}\}$ such that $\{c_{\ell},u\}\notin E(T)$. Consider the connected component of $T-c_{\ell}$ containing $u$, and notice that apart from $u$, it contains only vertices belonging to $\bigcup_{i\in[3n]}V_{i}$, none of which has an edge with $c_{\ell}$. In that case, $T$ is disconnected, which is a contradiction.
Next we show that for each $i\in[3n]$, there exists exactly one index $j\in[n]$ such that at least one vertex in $V_{i}$ is adjacent to $v_{j}$ in $T$. The proof is analogous to the one of Claim˜1.
Claim 3.
For all $i\in[3n]$, there exists $j\in[n]$ such that $N_{T}(V_{i})\cap Q=\{v_{j}\}$.
{claimproof}
[…]
$R\geq 2+\sum_{j\in[n]}(M+n-1)+\sum_{u\in\bigcup_{i\in[3n]}V_{i}}2(n-1)=2+n(M+n-1)+2(n-1)nB=2+nk>nk,$
which is a contradiction.
For $j\in[n]$, let $A_{j}=\{i\in[3n]\mid\exists u\in V_{i},\,\{u,v_{j}\}\in E(T)\}$. Due to Claim˜3 it follows that $(A_{1},\dots,A_{n})$ is a partition of $[3n]$. In particular, the set of vertices of the component of $T-\{c_{\ell},v_{j}\}$ containing $v_{j}$ is $\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. This implies that
$\mathrm{cng}_{G,T}(\{c_{\ell},v_{j}\})$ $=|E(\{v_{j}\},C)|+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[3n]\setminus A_{j}}V_{i})|+|E(\textstyle\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i},Q\setminus\{v_{j}\})|$
$=M+\sum_{i\in[3n]\setminus V_{j}}a_{i}+(n-1)+(n-1)\sum_{i\in V_{j}}a_{i}$
$=M+(n-1)+nB+(n-2)\sum_{i\in V_{j}}a_{i}.$
Combining this with the assumption $\mathrm{cng}_{G,T}(c_{\ell},v_{j})\leq k=M+(n-1)+2(n-1)B$, we obtain that $\sum_{i\in V_{j}}a_{i}\leq B$. (Recall that $n\geq 3$.) Since $\sum_{i\in[3n]}a_{i}=nB$, we have $\sum_{i\in V_{j}}a_{i}=B$ for all $j\in[n]$. ∎
By Lemmas˜3.15 and 3.16, LABEL:thm:hardness:modularwidth follows.
3.3 Maximum degree
The last result of Section˜3 is the NP-hardness on graphs of constant maximum degree. We first present an alternative gadget for the double-weighted edges which we subsequently use in our proof.
Double-weighted edges.
For the sake of simplicity in our reductions, we will use the concept of double-weighted edges introduced by Luu and Chrobak [algorithmica/LuuC24]. A double edge-weighted graph $G=(V,E;\mathrm{w}_{1},\mathrm{w}_{2})$ is a graph with two edge-weight functions $\mathrm{w}_{1},\mathrm{w}_{2}\colon E\to\mathbb{Z}^{+}$. For simplicity, let $\mathrm{w}(e)$ denote the tuple $(\mathrm{w}_{1}(e),\mathrm{w}_{2}(e))$ for $e\in E$. Let $T$ be a spanning tree of $G$. When considering the congestion of $T$, the double weights of the edges work slightly differently from the ordinal (single) edge weight considered in Section˜3.1. If $e\notin E(T)$, then it contributes $\mathrm{w}_{1}(e)$ to the congestions of the edges in the detour for $e$ in $T$; if $e\in E(T)$, it contributes $\mathrm{w}_{2}(e)$ to the congestion of itself. That is, for $e\in E(T)$, it holds that
$\mathrm{cng}_{G,T}(e)=\mathrm{w}_{1}(E(V(T_{e,1}),V(T_{e,2}))\setminus\{e\})+\mathrm{w}_{2}(e).$
Luu and Chrobak [algorithmica/LuuC24] showed that for every positive integer $k$, a double-weighted edge $e$ with $\mathrm{w}_{1}(e)\leq\mathrm{w}_{2}(e)<k$ can be replaced with a gadget consisting of unweighted edges (i.e., edges $e^{\prime}$ with $\mathrm{w}_{1}(e^{\prime})=\mathrm{w}_{2}(e^{\prime})=1$) without changing the property of having spanning tree congestion at most $k$. Their gadget increases the degree of one endpoint of $e$ by $\mathrm{w}_{1}(e)-1$ and the other by $\mathrm{w}_{1}(e)\cdot(\mathrm{w}_{2}(e)-\mathrm{w}_{1}(e)+1)-1$. Because of this increase of the degree, and as in the following reduction proving Theorem˜3.19, $w_{2}$ is unbounded, we cannot use their gadget in our proof.
In the following, we present an alternative gadget for double-weighted edges that does not increase the degree too much. For an integer $n\geq 2$, the $n\times n$ grid is the Cartesian product of two $n$-vertex paths. We call the degree-$2$ vertices in a grid its corners. It is known that the spanning tree congestion of the $n\times n$ grid is $n$ [dmgt/CastejonO09, dm/Hruska08].
Lemma 3.17.
lem:double-weighted_edge Let $k$ be a positive integer and $G=(V,E;\mathrm{w}_{1},\mathrm{w}_{2})$ be a double edge-weighted graph with an edge $e=\{u,v\}\in E$ satisfying that $\mathrm{w}_{1}(e)<\mathrm{w}_{2}(e)<k$. Let $G^{\prime}$ be the graph obtained from $G$ by the following modification (see Figure˜3 (left)):
  1. 1.
    remove $e$;
  2. 2.
    add $\mathrm{w}_{1}(e)$ copies of the $(\mathrm{w}_{2}(e)-\mathrm{w}_{1}(e)+1)\times(\mathrm{w}_{2}(e)-\mathrm{w}_{1}(e)+1)$ grid;
  3. 3.
    for each grid added in the previous step, add edges $\{u,c\}$ and $\{v,c^{\prime}\}$, where $c$ is an arbitrary corner of the grid and $c^{\prime}$ is the opposite corner (i.e., the corner furthest from $c$);
  4. 4.
    for each new edge $f\in E(G^{\prime})\setminus E(G)$, set $\mathrm{w}_{1}(f)=\mathrm{w}_{2}(f)=1$.
Then, $\mathrm{stc}(G)\leq k$ if and only if $\mathrm{stc}(G^{\prime})\leq k$. The degrees of $u$ and $v$ increase by $\mathrm{w}_{1}(e)-1$ and the maximum degree among newly added vertices is at most $4$.
Figure 3: (Left) The gadget for a double-weighted edge $e=\{u,v\}$ with $(\mathrm{w}_{1}(e),\mathrm{w}_{2}(e))=(3,5)$. There are $\mathrm{w}_{1}(e)$ (=3=3) grids connected to $\{u,v\}$ and each grid is of size $(\mathrm{w}_{2}(e)-\mathrm{w}_{1}(e)+1)\times(\mathrm{w}_{2}(e)-\mathrm{w}_{1}(e)+1)$ ($3\times 3$). (Center & Right) The intersection of the gadget and the spanning tree $T^{\prime}$ of $G^{\prime}$ obtained from a spanning tree $T$ of $G$ for the cases $e\in E(T)$ and $e\notin E(T)$: in the former case, the edges $\{v_{1},v\},\{v_{2},v\},\{v_{3},v\}$ and some edges in $\Gamma_{1}$ contribute to the congestion of the uuvv path in $T^{\prime}$; in the latter, only the edges $\{v_{1},v\},\{v_{2},v\},\{v_{3},v\}$ do.
Proof 3.18.
The claims about the vertex degrees hold clearly from the construction. In the following, we show that $\mathrm{stc}(G)\leq k$ if and only if $\mathrm{stc}(G^{\prime})\leq k$.
For simplicity, let $\mathrm{w}(e)=(a,b)$. We call the $a$ copies of the $(b-a+1)\times(b-a+1)$ grid $\Gamma_{1},\dots,\Gamma_{a}$. For $i\in[a]$, let $u_{i}$ and $v_{i}$ denote the neighbors of $u$ and $v$ in $\Gamma_{i}$, respectively.
[…]
*The only-if direction. Assume that $\mathrm{stc}(G)\leq k$ and $T$ is a spanning tree of $G$ with congestion at most $k$. We handle the cases $e\in E(T)$ and $e\notin E(T)$ separately. In both cases, let $T_{i}$ be a spanning tree of $\Gamma_{i}$ with congestion ba+1b-a+1. Such a spanning tree exists as the spanning tree congestion of the $n\times n$ grid is $n$ [dmgt/CastejonO09, dm/Hruska08].
We first consider the case where $e\in E(T)$. From $T$, we construct a spanning tree $T^{\prime}$ of $G^{\prime}$ by removing $e$, adding the edges in $T_{i}$ and the edge $\{u,u_{i}\}$ for $i\in[a]$, and adding the edge $\{v,v_{1}\}$ (see Figure˜3 (center)); that is,
$E(T^{\prime})=(E(T)\setminus\{e\})\cup\bigcup_{i\in[a]}(E(T_{i})\cup\{\{u,u_{i}\}\})\cup\{\{v,v_{1}\}\}.$
We can see that $T^{\prime}$ is a spanning tree of $G^{\prime}$. Observe that the edges in $E(G^{\prime})\setminus E(G)$ do not use the edges in $E(T)\setminus E(T^{\prime})$ in their detours in $T^{\prime}$. Observe also that for each edge $f\in E(G)\setminus\{e\}$, if the detour $P$ for $f$ in $T$ includes $e$, then the detour for $f$ in $T^{\prime}$ is obtained from $P$ by replacing $e$ with the uuvv path in $T^{\prime}$.
Let $f\in E(T^{\prime})$. If $f\in E(T)\setminus\{e\}$, then $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)=\mathrm{cng}_{G,T}(f)$. In the following, assume that $f\notin E(T)$. If $f=\{u,u_{i}\}$ for some $i\in[2,a]$, then $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)=|\{\{u,u_{i}\},\{v,v_{i}\}\}|=2\leq k$. If $f\in E(\Gamma_{i})$ for some $i\in[a]$ and $f$ does not belong to the uuvv path in $T^{\prime}$, then $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)\leq\mathrm{cng}_{\Gamma_{i},T_{i}}(f)+1\leq(b-a+1)+1\leq k$, where the term +1+1 may appear by the detour for $\{v,v_{i}\}$ when $i\neq 1$. In the remaining cases, $f$ belongs to the uuvv path in $T^{\prime}$. If $f$ is $\{u,u_{1}\}$ or $\{v,v_{1}\}$, then
$\mathrm{cng}_{G^{\prime},T^{\prime}}(f)=\mathrm{w}_{1}(E(V(T_{e,1}),V(T_{e,2}))\setminus\{e\})+|\{f\}\cup\{\{v,v_{i}\}\mid i\in[2,a]\}|.$
Similarly, if $f\in E(\Gamma_{1})$, then
$\mathrm{cng}_{G^{\prime},T^{\prime}}(f)\leq\mathrm{w}_{1}(E(V(T_{e,1}),V(T_{e,2}))\setminus\{e\})+|\{\{v,v_{i}\}\mid i\in[2,a]\}|+\mathrm{cng}_{\Gamma_{1},T_{1}}(f).$
In both cases, $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)\leq\mathrm{cng}_{G,T}(e)\leq k$ holds since $|\{\{v,v_{i}\}\mid i\in[2,a]\}|+\mathrm{cng}_{\Gamma_{1},T_{1}}(f)\leq(a-1)+(b-a+1)=b=\mathrm{w}_{2}(e)$.
Next, consider the case where $e\notin E(T)$. From $T$, we construct a spanning tree $T^{\prime}$ of $G^{\prime}$ by adding the edges in $T_{i}$ and the edge $\{u,u_{i}\}$ for $i\in[a]$ (see Figure˜3 (right)); that is,
$E(T^{\prime})=E(T)\cup\bigcup_{i\in[a]}(E(T_{i})\cup\{\{u,u_{i}\}\}).$
Clearly, $T^{\prime}$ is a spanning tree of $G^{\prime}$. The detours for the $a$ edges $\{v,v_{i}\}$ ($i\in[a]$) contain the uuvv path in $T$. No other edge in $E(T^{\prime})\setminus E(T)$ contributes to the congestions of edges in $E(T)$. Thus, if $f\in E(T)$ and $f$ does not belong to the uuvv path in $T$, then $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)=\mathrm{cng}_{G,T}(f)\leq k$. If $f\in E(T)$ and $f$ belongs to the uuvv path in $T$, then $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)=\mathrm{cng}_{G,T}(f)-\mathrm{w}_{1}(e)+|\{\{v,v_{i}\}\mid i\in[a]\}|=\mathrm{cng}_{G,T}(f)\leq k$. Finally, we consider the case where $f\in E(T^{\prime})\setminus E(T)$. If $f$ is $\{u,u_{i}\}$ for some $i\in[a]$, then $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)=|\{\{u,u_{i}\},\{v,v_{i}\}\}|=2\leq k$. Otherwise, $f\in E(\Gamma_{i})$ for some $i\in[a]$ and $\mathrm{cng}_{G^{\prime},T^{\prime}}(f)\leq\mathrm{cng}_{\Gamma_{i},T_{i}}(f)+1\leq(b-a+1)+1\leq b+1\leq k$, where the term +1+1 is necessary if $f$ is on the $u_{i}$$v_{i}$ path.
[…]
*The if direction. Assume that $\mathrm{stc}(G^{\prime})\leq k$ and $T^{\prime}$ is a spanning tree of $G^{\prime}$ with congestion at most $k$. Since the vertex set $\{u,v\}$ is a vertex-cut that separates the gadget replacing $e$ and the rest, the uuvv path $P$ in $T^{\prime}$ is either completely contained in the gadget or not intersecting it at all.
Let us first consider the case where the uuvv path $P$ is contained in the gadget. From $T^{\prime}$, we construct a spanning tree $T$ of $G$ by removing the gadget and then adding $e$; that is, $E(T)=(E(T^{\prime})\cap E(G))\cup\{e\}$.
For each $f\in E(G)$, if the detour for $f$ in $T^{\prime}$ contains $P$, then its detour in $T$ is obtained from the one in $T^{\prime}$ by replacing $P$ with $e$; otherwise, the detour for $f$ in $T$ is the same as the one in $T^{\prime}$. This implies that $\mathrm{cng}_{G,T}(f)=\mathrm{cng}_{G^{\prime},T^{\prime}}(f)$ for $f\in E(T)\setminus\{e\}$.
Now we show that $\mathrm{cng}_{G,T}(e)\leq k$. There is exactly one index $i\in[a]$ such that $P$ contains both $\{u,u_{i}\},\{v,v_{i}\}$ and a $u_{i}$$v_{i}$ path. It is known that in every spanning tree of the $n\times n$ grid, the path from a corner to the opposite corner contains an edge with congestion at least $n$ [dm/Hruska08]. This implies that there is an edge $e_{i}$ in $P$ such that at least ba+1b-a+1 edges in $\Gamma_{i}$ use $e_{i}$ in their detours in $T^{\prime}$. For each $j\in[a]\setminus\{i\}$, one of $\{u,u_{j}\}$ and $\{v,v_{j}\}$ uses $P$ (and thus $e_{i}$) in its detour in $T^{\prime}$. In total, at least $b$ edges in the gadget use $e_{i}$ in their detours. Furthermore, if an edge $f\in E(G)\setminus\{e\}$ uses $e$ in its detour in $T$, then $f$ uses $P$ (and thus $e_{i}$) in its detour in $T^{\prime}$. Thus we have
$\mathrm{cng}_{G,T}(e)$ $=\mathrm{w}_{1}(E(V(T_{e,1}),V(T_{e,2}))\setminus\{e\})+\mathrm{w}_{2}(e)$
$=\mathrm{w}_{1}(E(V(T_{e,1}),V(T_{e,2}))\setminus\{e\})+b\leq\mathrm{cng}_{G^{\prime},T^{\prime}}(e_{i})\leq k.$
Next we consider the case where the uuvv path $P$ does not intersect the gadget. From $T^{\prime}$, we construct a spanning tree $T$ of $G$ by removing the gadget; that is, $E(T)=E(T^{\prime})\cap E(G)$. If $f\in E(T)\setminus E(P)$, the edges using $f$ for their detours are the same in $T$ and $T^{\prime}$, and thus $\mathrm{cng}_{G,T}(f)=\mathrm{cng}_{G^{\prime},T^{\prime}}(f)\leq k$.
Now assume that $f\in E(P)$. If an edge in $G$ uses $f$ in its detour in $T$, then it uses $f$ in its detour in $T^{\prime}$ as well. For each $i\in[a]$, one of $\{u,u_{j}\}$ and $\{v,v_{j}\}$ uses $P$ (and thus $f$) in its detour in $T^{\prime}$. On the other hand, $e$ uses $P$ in its detour in $T$. Thus we have
$\mathrm{cng}_{G,T}(f)$ $=\mathrm{w}_{1}(E(V(T_{f,1}),V(T_{f,2}))\setminus\{e,f\})+\mathrm{w}_{1}(e)+\mathrm{w}_{2}(f)$
$=\mathrm{w}_{1}(E(V(T_{f,1}),V(T_{f,2}))\setminus\{e,f\})+a+\mathrm{w}_{2}(f)\leq\mathrm{cng}_{G^{\prime},T^{\prime}}(f)\leq k.$
The problem (3, B2)-SAT (also appearing in the literature as 2P2N-3SAT) is a restricted version of 3-SAT: an instance of (3, B2)-SAT consists of a set $X$ of $n$ variables and a set $C$ of $m$ clauses such that each clause has exactly three literals corresponding to three different variables and each variable appears exactly twice positively and exactly twice negatively. It is known that (3, B2)-SAT is NP-complete [eccc/ECCC-TR03-049], even if the formula contains only monotone clauses [dam/DarmannD21].
Theorem 3.19.
Spanning Tree Congestion is NP-complete on graphs of maximum degree at most $8$.
Construction.
Let $(X,C)$ be an instance of (3, B2)-SAT with $X=\{x_{1},\dots,x_{n}\}$ and $C=\{c_{1},\dots,c_{m}\}$. We assume that $m\geq 3$ (otherwise the problem becomes trivial). Set $k=2m+3$ ($\geq 9$). From $(X,C)$, we construct a double edge-weighted graph $G=(V,E;\mathrm{w}_{1},\mathrm{w}_{2})$ as follows (see Figure˜4).
  • For $i\in[n]$, take a cycle $(x_{i},y_{i},\bar{x}_{i},z_{i})$ of four new vertices.
    • Set $\mathrm{w}(\{x_{i},y_{i}\})=\mathrm{w}(\{\bar{x}_{i},y_{i}\})=(4,k-3)$ and $\mathrm{w}(\{x_{i},z_{i}\})=\mathrm{w}(\{\bar{x}_{i},z_{i}\})=(1,1)$.
  • For $i\in[n-1]$, add the edge $\{z_{i},z_{i+1}\}$, thus forming the path $(z_{1},\dots,z_{n})$.
    • Set $\mathrm{w}(\{z_{i},z_{i+1}\})=(3,3)$.
  • For $j\in[m]$, take a new vertex $c_{j}$.
  • For $i\in[n]$ and $j\in[m]$, add the edge $\{x_{i},c_{j}\}$ (resp. $\{\bar{x}_{i},c_{j}\}$) if $x_{i}\in c_{j}$ (resp. $\bar{x}_{i}\in c_{j}$).
    • Set $\mathrm{w}(\{x_{i},c_{j}\})=(1,k-2)$ (resp. $\mathrm{w}(\{\bar{x}_{i},c_{j}\})=(1,k-2)$) if the edge exists.
Figure 4: The construction of $G$, where $c_{1}=\{x_{1},\bar{x}_{2},x_{3}\}$, $c_{2}=\{x_{1},x_{2},\bar{x}_{3}\}$, $c_{3}=\{\bar{x}_{1},\bar{x}_{2},x_{3}\}$, and $c_{4}=\{\bar{x}_{1},x_{2},\bar{x}_{3}\}$.
To prove Theorem˜3.19, it suffices to prove the following two claims.
  1. 1.
    In polynomial time, we can construct an unweighted graph $G^{\prime}$ such that
    • $\mathrm{stc}(G)\leq k$ if and only if $\mathrm{stc}(G^{\prime})\leq k$;
    • the maximum degree of $G^{\prime}$ is at most $8$.
  2. 2.
    $\mathrm{stc}(G)\leq k$ if and only if $(X,C)$ is a yes-instance of (3, B2)-SAT.
LABEL:lem:maxdeg_unweighting shows the first claim and LABEL:lem:maxdeg_equivalence shows the second one.
Lemma 3.20.
lem:maxdeg_unweighting In time polynomial in $m+n$, one can construct an unweighted graph $G^{\prime}=(V^{\prime},E^{\prime})$ from the double edge-weighted graph $G=(V,E;\mathrm{w}_{1},\mathrm{w}_{2})$ with maximum degree at most $8$ such that $\mathrm{stc}(G)\leq k$ if and only if $\mathrm{stc}(G^{\prime})\leq k$.
Proof 3.21.
We keep the edges with weight $(1,1)$. For each edge $\{u,v\}$ with weight $(3,3)$, we use Proposition˜3.4 and replace the edge $\{u,v\}$ with three internally vertex disjoint uuvv paths of length $2$. For each edge with weight (4,k3)(4,k-3) or (1,k2)(1,k-2), we replace the edge with the gadget in LABEL:lem:double-weighted_edge. We call the obtained unweighted graph $G^{\prime}=(V^{\prime},E^{\prime})$. Since $k=2m+3$, the construction can be done in time polynomial in $m+n$. By Propositions˜3.4 and LABEL:lem:double-weighted_edge, $\mathrm{stc}(G)\leq k$ if and only if $\mathrm{stc}(G^{\prime})\leq k$.
Now we show that $G^{\prime}$ has maximum degree at most $8$. Observe that $V^{\prime}\supseteq V$ and that the maximum degree of the vertices in $V^{\prime}\setminus V$ is at most $4$. For each $i\in[n]$, $\deg_{G^{\prime}}(z_{i})\leq\deg_{G}(z_{i})+4\leq 8$, $\deg_{G^{\prime}}(y_{i})=\deg_{G}(y_{i})+6=8$, $\deg_{G^{\prime}}(x_{i})=\deg_{G}(x_{i})+3=7$, and $\deg_{G^{\prime}}(\bar{x}_{i})=\deg_{G}(\bar{x}_{i})+3=7$. For each $j\in[m]$, $\deg_{G^{\prime}}(c_{j})=\deg_{G}(c_{j})=3$.
Lemma 3.22.
lem:maxdeg_equivalence $\mathrm{stc}(G)\leq k$ if and only if $(X,C)$ is a yes-instance of (3, B2)-SAT.
Proof 3.23.
To show the if direction, assume that $(X,C)$ is a yes-instance of (3, B2)-SAT and $\alpha\colon X\to\{0,1\}$ is a satisfying assignment of $(X,C)$. From $\alpha$, we construct a spanning tree $T$ of $G$ as follows (see Figure˜5).
  • Take all edges in the path $(z_{1},\dots,z_{n})$.
  • For $i\in[n]$, take both edges incident to $y_{i}$.
  • For $i\in[n]$, if $\alpha(x_{i})=1$, then take the edge $\{x_{i},z_{i}\}$; otherwise, take the edge $\{\bar{x}_{i},z_{i}\}$;
  • For $j\in[m]$, fix one literal $\ell_{i}\in c_{j}$ evaluated to $1$ under $\alpha$ (that is, if $\ell_{i}=x_{i}$, then $\alpha(x_{i})=1$; otherwise $\alpha(x_{i})=0$) and take the edge $\{c_{j},\ell_{i}\}$.
We first show that $\mathrm{cng}_{G,T}(\{z_{i},z_{i+1}\})\leq k$ for $i\in[n-1]$. Observe that an edge $e\in E(G)\setminus E(T)$ uses $\{z_{i},z_{i+1}\}$ in its detour only if $e$ is incident to some $c_{j}$. In total, there are exactly $3m$ edges incident to some $c_{j}$ and $m$ of them are included in $T$. Therefore, $\mathrm{cng}_{G,T}(\{z_{i},z_{i+1}\})\leq 2m+\mathrm{w}_{2}(\{z_{i},z_{i+1}\})=2m+3=k$.
We next show that, for $i\in[n]$, each edge incident to $y_{i}$ has congestion $k$. By symmetry, assume that $\{x_{i},z_{i}\}\in E(T)$. Let $c_{j_{1}},c_{j_{2}}$ be the clauses that $\bar{x}_{i}$ appears in. Then we have $\mathrm{cng}_{G,T}(\{\bar{x}_{i},y_{i}\})=\mathrm{w}_{1}(\{\{\bar{x}_{i},z_{i}\},\{\bar{x}_{i},c_{j_{1}}\},\{\bar{x}_{i},c_{j_{2}}\}\})+\mathrm{w}_{2}(\{\bar{x}_{i},y_{i}\})=3+(k-3)=k$. Similarly, $\mathrm{cng}_{G,T}(\{x_{i},y_{i}\})=\mathrm{w}_{1}(\{\{\bar{x}_{i},z_{i}\},\{\bar{x}_{i},c_{j_{1}}\},\{\bar{x}_{i},c_{j_{2}}\}\})+\mathrm{w}_{2}(\{x_{i},y_{i}\})=3+(k-3)=k$.
Now we show that, for $j\in[m]$, the edge incident to $c_{j}$ has congestion $k$. Let $c_{j}=\{\ell_{i_{1}},\ell_{i_{2}},\ell_{i_{3}}\}$, where $\{c_{j},\ell_{i_{1}}\}\in E(T)$. Since $c_{j}$ is a leaf of $T$, it holds that $\mathrm{cng}_{G,T}(\{c_{j},\ell_{i_{1}}\})=\mathrm{w}_{1}(\{c_{j},\ell_{i_{2}}\},\{c_{j},\ell_{i_{3}}\})+\mathrm{w}_{2}(\{c_{j},\ell_{i_{1}}\})=2+(k-2)=k$.
Finally, we show that, for $i\in[n]$, the edge in $T$ between $z_{i}$ and $x_{i}$ or $\bar{x}_{i}$ has congestion at most 88 ($\leq k$). Assume by symmetry that $\{z_{i},x_{i}\}\in E(T)$. Let $T_{x_{i}}$ be the connected component of $T-\{z_{i},x_{i}\}$ including $x_{i}$. The vertex set $V(T_{x_{i}})$ includes $x_{i},y_{i},\bar{x}_{i}$ and zero, one, or two vertices in $\{c_{1},\dots,c_{m}\}$ adjacent to $x_{i}$. Thus, we have $|E(V(T_{x_{i}}),V(G)\setminus V(T_{x_{i}}))|\leq 8$. Since every edge $e$ in $E(V(T_{x_{i}}),V(G)\setminus V(T_{x_{i}}))$ satisfies $\mathrm{w}_{1}(e)=1$, it holds that $\mathrm{cng}_{G,T}(\{z_{i},x_{i}\})=\mathrm{w}_{1}(E(V(T_{x_{i}}),V(G)\setminus V(T_{x_{i}}))\setminus\{\{z_{i},x_{i}\}\})+\mathrm{w}_{2}(\{z_{i},x_{i}\})\leq 8$.
Figure 5: The construction of $T$, where $\alpha=\{(x_{1},1),(x_{2},0),(x_{3},0)\}$.
To show the only-if direction, assume that $\mathrm{stc}(G)\leq k$ and $T$ is a spanning tree of $G$ with congestion at most $k$. We first show the following series of claims.
Claim 4.
For $i\in[n]$, $\{x_{i},y_{i}\},\{\bar{x}_{i},y_{i}\}\in E(T)$.
{claimproof}
Since $\deg_{G}(y_{i})=2$, at least one of $\{x_{i},y_{i}\}$ and $\{\bar{x}_{i},y_{i}\}$ belongs to $T$. Suppose to the contrary that one of them does not belong to $T$. By symmetry, we may assume that $\{x_{i},y_{i}\}\notin E(T)$, and thus $\{\bar{x}_{i},y_{i}\}\in E(T)$. Then, we have $\mathrm{cng}_{G,T}(\{\bar{x}_{i},y_{i}\})=\mathrm{w}_{1}(\{x_{i},y_{i}\})+\mathrm{w}_{2}(\{\bar{x}_{i},y_{i}\})=4+k-3>k$, a contradiction.
Claim 5.
For $i\in[n-1]$, $\{z_{i},z_{i+1}\}\in E(T)$.
{claimproof}
Suppose to the contrary that $\{z_{i},z_{i+1}\}\notin E(T)$ for some $i\in[n-1]$. Let $P$ be the $z_{i}$$z_{i+1}$ path in $T$. Observe that $G-\{c_{1},\dots,c_{m}\}$ is a cactus graph that consists of the path $(z_{1},\dots,z_{n})$ and $4$-cycles attached to the vertices on the path. Therefore, the edge $\{z_{i},z_{i+1}\}$ is the unique $z_{i}$$z_{i+1}$ path in $G-\{c_{1},\dots,c_{m}\}$. This implies that $P$ contains $c_{j}$ for some $j\in[m]$. Let $e\in E(P)$ be an edge incident to $c_{j}$. Then, we have $\mathrm{cng}_{G,T}(e)\geq\mathrm{w}_{1}(\{z_{i},z_{i+1}\})+\mathrm{w}_{2}(e)=3+k-2>k$, a contradiction.
Claim 6.
For $i\in[n]$, exactly one of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ belongs to $T$.
{claimproof}
By Claim˜4, at most one of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ may belong to $T$. Suppose to the contrary that none of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ belong to $T$. Let $\{y_{i},\ell_{i}\}$ and $\{\ell_{i},c_{j}\}$ be the first and second edges, respectively, in the $y_{i}$$z_{i}$ path in $T$, where $\ell_{i}\in\{x_{i},\bar{x}_{i}\}$. Observe that the edge set $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$ is not an edge-cut in $G$: the graph $G-\{x_{i},y_{i},\bar{x}_{i}\}$ is connected; and there is an edge not included in $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$ that connects $G-\{x_{i},y_{i},\bar{x}_{i}\}$ and $G[\{x_{i},y_{i},\bar{x}_{i}\}]$ (e.g., there is an edge between $x_{i}$ and $c_{j^{\prime}}$ for some $j^{\prime}\neq j$). Therefore, the edge set $E(V(T_{\{x_{i},c_{j}\},1}),V(T_{\{x_{i},c_{j}\},2}))$ is a proper super set of $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$. In that case,
$\mathrm{cng}_{G,T}(\{x_{i},c_{j}\})$ $>\mathrm{w}_{1}(\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\}\})+\mathrm{w}_{2}(\{x_{i},c_{j}\})=2+(k-2)=k,$
which is a contradiction.
By the claims above, $T-\{c_{1},\dots,c_{m}\}$ is connected. Since $\{c_{1},\dots,c_{m}\}$ is an independent set, we can conclude that each $c_{j}$ is a leaf of $T$.
Claim 7.
For $j\in[m]$, $c_{j}$ is a leaf of $T$.
Claim 8.
Let $i\in[n]$ and $\ell_{i}\in\{x_{i},\bar{x}_{i}\}$. If $\{\ell_{i},c_{j}\}\in E(T)$ for some $j\in[m]$, then $\{\ell_{i},z_{i}\}\in E(T)$.
{claimproof}
By symmetry, it suffices to show that if $\{x_{i},c_{j}\}\in E(T)$ for some $j\in[m]$, then $\{x_{i},z_{i}\}\in E(T)$. Suppose to the contrary that $\{x_{i},c_{j}\}\in E(T)$ for some $j\in[m]$ but $\{x_{i},z_{i}\}\notin E(T)$. By Claim˜4, $\{x_{i},y_{i}\},\{\bar{x}_{i},y_{i}\}\in E(T)$. By Claim˜6, $\{\bar{x}_{i},z_{i}\}\in E(T)$. Observe that the edge $\{x_{i},z_{i}\}$ and the edges incident to $c_{j}$ except for $\{x_{i},c_{j}\}$ use $\{x_{i},y_{i}\}$ in their detours in $T$. Let $j^{\prime}\in[m]$ be the index such that $j^{\prime}\neq j$ and $\{x_{i},c_{j^{\prime}}\}\in E(G)$. If $\{x_{i},c_{j^{\prime}}\}\notin E(T)$, then $\{x_{i},c_{j^{\prime}}\}$ uses $\{x_{i},y_{i}\}$ in its detour in $T$. Otherwise, the edges incident to $c_{j^{\prime}}$ except for $\{x_{i},c_{j^{\prime}}\}$ use $\{x_{i},y_{i}\}$ in their detours in $T$. In total, at least four edges not including $\{x_{i},y_{i}\}$ itself use $\{x_{i},y_{i}\}$ in their detours in $T$. Therefore, we have $\mathrm{cng}_{G,T}(\{x_{i},y_{i}\})\geq 4+\mathrm{w}_{2}(\{x_{i},y_{i}\})=4+(k-3)>k$, a contradiction.
Now we construct an assignment $\alpha\colon X\to\{0,1\}$ from $T$ by setting $\alpha(x_{i})=1$ if $\{x_{i},z_{i}\}\in E(T)$ and $\alpha(x_{i})=0$ if $\{\bar{x}_{i},z_{i}\}\in E(T)$. Note that, by Claim˜6, exactly one of these cases happens for each $i\in[n]$. We show that $\alpha$ is a satisfying assignment of $(X,C)$. Let $j\in[m]$. There is an index $i\in[n]$ such that $x_{i}$ or $\bar{x}_{i}$ is adjacent to $c_{j}$ in $T$. By symmetry, assume that $\{x_{i},c_{j}\}\in E(T)$. By Claim˜8, $\{x_{i},z_{i}\}\in E(T)$ holds. Therefore, we have $\alpha(x_{i})=1$, and thus the clause $c_{j}$ is satisfied by $\alpha$.
4 Algorithms for Bounded Treewidth
In this section we take a second look at the complexity of Spanning Tree Congestion parameterized by treewidth, a problem shown to be W[1]-hard in Corollary˜3.2. One way to deal with this hardness is to consider additional parameters, so we begin by presenting in Section˜4.1 an FPT algorithm parameterized by treewidth plus the desired congestion $k$. Our algorithm follows the standard technique of performing dynamic programming over a tree decomposition, though with a few necessary tweaks (informally, we have to guess the general structure of the spanning tree, including parts of the graph that will appear “in the future”).
We note here that the fact that Spanning Tree Congestion is FPT for this parameterization was already shown in [algorithmica/BodlaenderFGOL12], where it was proved that if $k$ is a parameter, then the problem is MSO2-expressible, hence solvable via Courcelle’s theorem. Nevertheless, we are still motivated to provide an explicit algorithm for the parameter “treewidth plus kk” for several reasons. First, using Courcelle’s theorem does not provide any usable upper bound on the running time, while we show our algorithm to run in ${(k+w)}^{{\cal O}(w)}n^{{\cal O}(1)}$, where $w$ is the treewidth; this implies, for instance, that Spanning Tree Congestion is in XP parameterized by treewidth alone (as $\mathrm{stc}(G)\leq n^{2}$ for all $G$), a fact that cannot be inferred using Courcelle’s theorem.
Second, and more importantly, having an explicit algorithm at hand, we are able to obtain an answer to the following natural question: given that solving Spanning Tree Congestion is hard parameterized by treewidth, is there an FPT algorithm that closely approximates the optimal congestion? By applying a technique of Lampis [icalp/Lampis14] which modifies exact DP algorithms to obtain approximate ones, we get an FPT approximation scheme, which runs in time ${(\frac{w}{\varepsilon})}^{{\cal O}(w)}n^{{\cal O}(1)}$ (that is, FPT in $w+\frac{1}{\varepsilon}$) and returns a $(1+\varepsilon)$-approximate solution, for any desired $\varepsilon>0$. This result naturally complements the problem’s hardness for treewidth and is presented in Section˜4.2. We complete that section by presenting a simple Win/Win argument which extends our algorithm to an algorithm that is FPT parameterized by clique-width plus $k$; this is based on a result of Gurski and Wanke [wg/GurskiW00] stating that graphs of bounded clique-width with no large complete bipartite subgraphs actually have bounded treewidth.
4.1 FPT Algorithm Parameterized by Treewidth and Congestion
In this section, we prove the following theorem:
Theorem 4.24.
Let $G$ be a graph with treewidth $w$, and let $k>0$. There is an algorithm that finds a spanning tree of $G$ with congestion $k$, if it exists, and runs in time ${(w+k)}^{{\cal O}(w)}n^{{\cal O}(1)}$.
Before we proceed with the formal proof of the theorem, we will give an intuition of the algorithm. Following the usual structure for bounded-treewidth graphs, we will use dynamic programming to compute solutions to the subgraphs corresponding to subtrees of the tree decomposition $\mathcal{T}$. For each such subgraph, we consider different possibilities for the solution to manifest on the root bag of the subtree; these different possibilities, which we call states, represent equivalent classes of solutions that are “compatible” with respect to the rest of the graph, and thus it is sufficient to compute a feasible solution for each such class.
To obtain the algorithm, we simply need to specify the states, as well as how to recursively obtain feasible solutions for each. Our states are composed of a tree, which we call the skeleton, as well as values of congestion for each of its edges. The skeleton of a solution $T$ for a given bag $X_{t}$ intuitively represents how $T$ connects the vertices of $X_{t}$, and it can be obtained by contracting the vertices in $V\setminus X_{t}$ that have degree 1 or 2 in $T$. Thus, the skeleton is a tree containing the vertices in $X_{t}$, plus at most $|X_{t}|$ vertices with degree at least $3$. The congestion values correspond to the congestion induced on an edge by edges in $\mathcal{T}_{t}$, the subtree of $\mathcal{T}$ rooted at $t$.
There is one further particularity that we must consider: when constructing a skeleton from a solution, some of the resulting edges may represent paths in the subtree rooted at $t$, while others represent paths outside of this subtree. Thus, we label each edge of the skeleton with one of three types: a present edge is simply an edge between two vertices in $X_{t}$; a past edge represents a path in $\mathcal{T}_{t}$; and a future edge represents a path outside $\mathcal{T}_{t}$, that is, one that is not (yet) in the solution, but that must be added to make it compatible with the state. We similarly label the vertices with the type present if they are in $X_{t}$, past if they are not in $X_{t}$ but in $\mathcal{T}_{t}$, and future otherwise.
Throughout the section, we assume familiarity with the definition and usual notation for treewidth (see e.g. [books/CyganFKLMPPS15, Chapter 7]). When referring to subgraphs of $G$, we often refer to subgraphs induced by subtrees of $\mathcal{T}$: we denote by $\mathcal{T}_{t}$ the subtree of $\mathcal{T}$ rooted at $t$, and by $G[S]$ the subgraph of $G$ induced by the vertices contained in bags of $S$, i.e. the subgraph $G\big[\bigcup_{t\in S}X_{t}\big]$; for convenience of notation, we write $G[\mathcal{T}_{t}]$ instead of $G[V(\mathcal{T}_{t})]$.
The algorithm starts by computing a nice tree decomposition $(\mathcal{T},\mathcal{X})$ for $G$ with width at most $2w+1$ and at most ${\cal O}(nw)$ bags, which can be computed in time $2^{{\cal O}(w)}\cdot n$ [Korhonen21].
We now formalize the definition of skeleton:
Definition 4.25.
Given a graph $X$, a skeleton $(S,\ell)$ is a tree $S$ together with a labeling $\ell\colon E(S)\cup V(S)\to\{-1,0,1\}$, such that:
  1. 1.
    $V(X)\subseteq V(S)$;
  2. 2.
    for any $v\in V(S)\setminus V(X)$, $\deg_{S}(v)\geq 3$;
  3. 3.
    for any $v\in V(S)$, $\ell(v)=0$ if and only if $v\in X_{t}$;
  4. 4.
    for any $uv\in E(S)$, $\ell(u)=\ell(uv)$ or $\ell(u)=0$ (similarly for $v$).
  5. 5.
    for any $uv\in E(S)$, if $\ell(uv)=0$ then $uv\in E(X)$;
If $(S,\ell)$ satisfies every property except Property 2, we call it a quasi-skeleton.
We denote by $\ell^{-1}(i)$, $i\in\{-1,0,1\}$ the graph obtained from the edges with label $i$ and their endpoints (which can have label $i$ or 0).
The labeling $\ell$ corresponds to the edge and vertex types and is encoded as 1-1 for past, 0 for present and $1$ for future.
We define a dynamic programming table $D$ with entries for every node $t\in\mathcal{T}$ and every triple $(S,\ell,c)$, where $(S,\ell)$ is a skeleton for $G[X_{t}]$ and $c\colon E(S)\to[0,k]$ is a function assigning a congestion to each edge. Informally, $D[t,(S,\ell,c)]$ represents a forest $F\subseteq G[\mathcal{T}_{t}]$ such that $F\cup\ell^{-1}(1)$ is a solution of congestion at most $k$ for $G[\mathcal{T}_{t}]$, and the congestion $c(uv)$ on an edge of the skeleton corresponds to the maximum congestion of the $u$-$v$-path in $F$. We formalize the desired properties by the definition below:
Definition 4.26.
We say that a forest $F\subseteq G[\mathcal{T}_{t}]$ is a consistent solution for $(t,(S,\ell,c))$ if:
  1. 1.
    $F$ is a forest of $G[\mathcal{T}_{t}]$ on the same vertex set;
  2. 2.
    For $uv\in E(S)$, if $\ell(uv)=0$, then $uv\in F$;
  3. 3.
    For $uv\in E(S)$, if $\ell(uv)=-1$, then $F$ contains a $u$-$v$-path $F_{uv}$ with edges from $G[\mathcal{T}_{t}]-E(X_{t})$};
  4. 4.
    $T^{\prime}:=F\cup\ell^{-1}(1)$ is a tree;
  5. 5.
    For any edge $e\in E(S)$ with $\ell(e)\in\{0,1\}$, the congestion in $T^{\prime}$ induced by $G[\mathcal{T}_{t}]-E(X_{t})$ on $e$ is $c(e)$;
  6. 6.
    For any edge $e\in E(S)$ with $\ell(e)=-1$, the congestion in $T^{\prime}$ induced by $G[\mathcal{T}_{t}]-E(X_{t})$ on every edge of $F_{e}$ is at most $c(e)$;
  7. 7.
    For any $e\in E(T^{\prime})$, the congestion in $T^{\prime}$ induced by $G[\mathcal{T}_{t}]-E(X_{t})$ on $e$ is at most $k$.
Before we describe the recursive rules of our algorithm, we show an operation called simplification which takes a quasi-skeleton and transforms it into a skeleton, which is necessary to keep the number of skeletons small enough.
Lemma 4.27.
lem:tw-alg:quasi-skel Let $X$ be a graph, $(S,\ell)$ be a quasi-skeleton for that graph, and $c\colon E(S)\to[k]$. We can obtain a skeleton $(S^{\prime},\ell^{\prime})$ for $X$ by starting with $(S^{\prime},\ell^{\prime})=(S,\ell)$, $c^{\prime}=c$ and iteratively contracting a vertex $v\in V(S^{\prime})\setminus V(X)$ with degree less than $3$ as follows:
  • if it has degree 1, we simply remove $v$ and its incident edge from $S^{\prime}$, $\ell^{\prime}$, $c$;
  • if it has degree 2 and label $\lambda=\ell(v)$, we replace $v$ and its incident edges by an edge $uw$ between its neighbors $u$ and $w$, and set $\ell^{\prime}(uw)=\lambda$, and adjust $c^{\prime}(uw)=\max\{c(uv),c(vw)\}$.
We name this process simplification.
Proof 4.28.
When the process completes, $(S^{\prime},\ell^{\prime})$ must satisfy Property 2, as we iteratively remove vertices that do not satisfy it.
For the remaining properties, we will show that an application of the lemma for a single step preserves them, and thus by repeated application the lemma holds. Properties 1 and 3 are not affected by removing vertices, and if an edge $uw$ is created, it has label $\ell(uw)=\ell(v)\in\{-1,1\}$, thus Property 5 is satisfied. Finally, since $\ell(v)\neq 0$, $\ell(uv)=\ell(vw)=\ell(v)$, and thus Property 4 is satisfied by setting $\ell(uw)=\ell(v)$.
We will now see how to construct the entries $D[t,(S,\ell,c)]$ recursively:
  • Leaf $t$: the only allowed skeleton is $S=(\varnothing,\varnothing)$, and $D[t,(S,\varnothing,\varnothing)]=(\varnothing,\varnothing)$.
  • Forget node $t$: let $t^{\prime}$ be the child of $t$ and let $\{v\}=X_{t^{\prime}}\setminus X_{t}$. For any solution $D[t^{\prime},(S^{\prime},\ell^{\prime},c^{\prime})]$, if $v$ has an incident edge with label $1$, the solution is invalid; otherwise, we set $D[t,(S,\ell,c)]=D[t^{\prime},(S^{\prime},\ell^{\prime},c^{\prime})]$ for $(S,\ell,c)$ obtained as follows:
    1. 1.
      we start with $(S,\ell,c)=(S^{\prime},\ell^{\prime},c^{\prime})$;
    2. 2.
      we add the congestion corresponding to the forgotten vertex $v$: for any edge $uv\in E(G)$, $u\in X_{t}$, increment $c(e)$ on each edge $e$ on the $u$-$v$-path in $S$; if $c(e)>k$ for any edge, the solution is invalid and the process stops;
    3. 3.
      then we mark $v$ as a past vertex ($\ell(v)=-1$) and apply simplification.
  • Introduce node $t$: let $t^{\prime}$ be the child of $t$ and let $\{v\}=X_{t}\setminus X_{t^{\prime}}$. For any $(S,\ell,c)$, we obtain $(S^{\prime},\ell^{\prime},c^{\prime})$ from $(S,\ell,c)$ by setting $\ell^{\prime}(v)=1$ and applying simplification. If $D[t^{\prime},(S^{\prime},\ell^{\prime},c^{\prime})]$ exists, we construct $D[t,(S,\ell,c)]$ from $D[t^{\prime},(S^{\prime},\ell^{\prime},c^{\prime})]$ by adding the vertex $v$ and every of its incident edges $e\in E(S)$ with label $\ell(e)=0$.
  • Join node $t$: given solutions $D[t_{1},(S,\ell_{1},c_{1})]$, $D[t_{2},(S,\ell_{2},c_{2})]$, a valid solution can be obtained if $c_{1}(e)+c_{2}(e)\leq k$, if $\ell_{1}(x)=-1$ implies $\ell_{2}(x)=1$ and vice-versa ($x\in V(S)\cup E(S)$), and $\ell_{1}(e)=0$ if and only if $\ell_{2}(e)=0$.
    We define $\ell(x)=\min\{\ell_{1}(x),\ell_{2}(x)\}$, $x\in V(S)\cup E(S)$} and $c(e)=c_{1}(e)+c_{2}(e)$, $e\in E(S)$, and set $D[t,(S,\ell,c)]=D[t_{1},(S,\ell_{1},c_{1})]\cup D[t_{2},(S,\ell_{2},c_{2})]$.
To obtain a solution to the problem, we simply compute $T=D[r,(\varnothing,\varnothing,\varnothing)]$. If $T$ is a consistent solution, then it is a forest containing the vertices $V(G)$ by Property 1, it is a tree by Property 4, and has congestion at most $k$ by Property 7. We therefore conclude that $T$ is a spanning tree of $G$ with congestion at most $k$, as desired.
LABEL:lem:tw-alg:correctn, LABEL:lem:tw-alg:completen and LABEL:lem:tw-alg:runtime show that $T$ is indeed a consistent solution, that if there is a feasible solution then our algorithm can always find it, and that the algorithm runs in the stated running time.
Lemma 4.29.
lem:tw-alg:correctn For any $t$ and $(S,\ell,c)$, if $D[t,(S,\ell,c)]$ exists, then it is a consistent solution.
Proof 4.30.
If $t\in V(\mathcal{T})$ is a leaf, then the properties of Definition˜4.26 hold trivially, as $G[\mathcal{T}_{t}]$ is an empty graph. We will now show by induction (from the leaves up) that the claim holds for every node $t$. Let $t\in V(\mathcal{T})$ be any node, and $(S,\ell,c)$ such that $F:=D[t,(S,\ell,c)]$ exists.
If $t$ is an introduce node, then $F=D[t^{\prime},(S^{\prime},\ell^{\prime},c^{\prime})]\cup\ell^{-1}(0)$ for some $(S^{\prime},\ell^{\prime},c^{\prime})$. In the dynamic program, $(S^{\prime},\ell^{\prime},c^{\prime})$ is obtained from $(S,\ell,c)$ by setting $\ell^{\prime}(v)=1$ and applying simplification. In this case, simplification means that if $v$ has degree $1$, it gets contracted, and if $v$ (or its parent in the previous case) have degree $2$, they get contracted. From the opposite perspective, we obtain $(S,\ell,c)$ from $(S^{\prime},\ell^{\prime},c^{\prime})$ by optionally subdividing an edge, and then either identifying $v$ as one of the vertices with label $1$ or adding it as a leaf child of a vertex with label 0 or $1$. The properties are preserved by subdivision since we assign the same congestion to both new edges as in the old one, and since edges have label $1$ by default and only have label 0 if they exist in $E(X_{t})$. Similarly, we can see that adding $v$ as a leaf does not affect the properties.
Furthermore, $G[\mathcal{T}_{t}]-E(X_{t})$ is equal to $G[\mathcal{T}_{t^{\prime}}]-E(X_{t^{\prime}})$ with an additional isolated vertex $v$, since all of the edges incident on $v$ in $G[\mathcal{T}_{t}]$ are also in $E(X_{t})$. Thus, as we only count congestion of edges in $G[\mathcal{T}_{t}]-E(X_{t})$, congestion does not change.
If $t$ is a forget node, we first show that adding the congestion does not change the properties. For every new edge $uv\in E(X_{t^{\prime}})$ incident on $v$, the construction increments $c(e)$ for every edge $e$ on the $u$-$v$-path in $S$; this precisely matches the edges of $T^{\prime}$ whose congestion is increased by the edge $uv$, as edges with label 0 and $1$ are present in $T^{\prime}$, and edges with label 1-1 are present in the form of a path. Property 7 is preserved since we check that no edge exceeds the congestion value of $k$.
The contraction operations used in simplification do not change $F$ and do not add any edges with label 0 or $1$, but affect Properties 3, 6, in particular when contracting a vertex $v^{\prime}$ with degree 2. However, the action of contracting $v^{\prime}$ simply concatenates the paths $F_{uv^{\prime}}$ and $F_{v^{\prime}w}$ to obtain the path $F_{uw}$, and setting $c(uw)=\max\{c(uv^{\prime}),c(v^{\prime}w)\}$ ensures that each edge in $F_{uv^{\prime}}$ has capacity at most $c(uw)$.
If $t$ is a join node, $F$ is obtained as the union of solutions $F_{1}:=D[t_{1},(S,\ell_{1},c_{1})]$, $F_{2}:=D[t_{2},(S,\ell_{2},c_{2})]$ for the children nodes $t_{1}$, $t_{2}$, and we assume that for every $e\in E(S)$, $x\in V(S)\cup E(S)$, $c_{1}(e)+c_{2}(e)\leq k$, that if $\ell_{1}(x)=-1$ then $\ell_{2}(x)=1$ and vice-versa, and that $\ell_{1}(e)=0$ if and only if $\ell_{2}(e)=0$. Properties 1, 2, 3 follow by definition of $F$ and of $\ell$ as $\ell(x)=\min\{\ell_{1}(x),\ell_{2}(x)\}$. In particular, the definition of $\ell$ implies that for every edge $uw\in E(S)$ such that $\ell(uw)=-1$, either $\ell_{1}(uw)=-1$ or $\ell_{2}(uw)=-1$, and thus Property 3 implies that a $u$-$w$-path exists in either $F_{1}$ or $F_{2}$, and thus in $F$. We can similarly see that $T=F\cup\ell^{-1}(1)$ is a tree, as we are equivalently taking $S$ and replacing each $uw\in E(S)$ with label 1-1 by a subtree that connects $u$ and $w$; as we know that $F_{1}\cup\ell_{1}^{-1}(1)$ and $F_{2}\cup\ell_{2}^{-1}(1)$ are both trees by induction, and that $\ell_{1}^{-1}$ and $\ell_{2}^{-1}$ are edge-disjoint (as they use edges from disjoint subtrees of $\mathcal{T}$), then $T$ must be acyclic, and thus a tree. Finally, we can see that each edge in $G[\mathcal{T}_{t}]-E(X_{t})$ must belong to either $G[\mathcal{T}_{t_{1}}]-E(X_{t})$ or $G[\mathcal{T}_{t_{2}}]-E(X_{t})$, and thus the corresponding congestion is accounted for in $S_{1}$ or $S_{2}$, respectively; by summing these congestion values, we ensure that Properties 5, 6, 7 are satisfied.
Lemma 4.31.
lem:tw-alg:completen Let $T$ be a tree with congestion at most $k$. Then for any $t$, there is $(S_{t},\ell_{t},c_{t})$ such that $D[t,(S,\ell,c)]$ exists.
Proof 4.32.
We will start by defining, for each $t\in V(\mathcal{T})$, a solution $F_{t}$ and a state $(S_{t},\ell_{t},c_{t})$, and then show that $F_{t}$ is consistent solution for $D[t,(S_{t},\ell_{t},c_{t})]$, and thus $D[t,(S_{t},\ell_{t},c_{t})]$ exists.
The solution $F_{t}$ can be obtained by restricting $T$ to the corresponding subgraph, that is, $F_{t}:=T\cap G[\mathcal{T}_{t}]$. $S_{t}$ can be obtained by labeling the vertices of $T$ according to $\mathcal{T}$, and then applying simplification. Concretely, we set $\ell(v)=0$, $v\in X_{t}$ and $\ell(e)=0$, $e\in T\cap E(X_{t})$; then, we set $\ell(v)=-1$ for $v\in V(G[\mathcal{T}_{t}])\setminus X_{t}$ and $\ell(v)=1$ if $v\in V\setminus V(G[\mathcal{T}_{t}])$; finally, an edge $e\in T\setminus E(X_{t})$ gets the non-zero label of one of its endpoints. For every edge $e\in E(S_{t})$, we take $c(e)$ to be the congestion induced by the edges in $G^{\prime}_{t}:={G[\mathcal{T}_{t}]-E(X_{t})}$ on $e$. We can now see that $(T,\ell_{t})$ is a quasi-skeleton for $G[X_{t}]$, and thus, by LABEL:lem:tw-alg:quasi-skel, we obtain the skeleton $(S_{t},\ell_{t})$ by simplification.
Next we show that $F_{t}$ is a consistent solution: by definition, it is a forest containing every vertex in $G^{\prime}_{t}$, every edge with label 0, and a path connecting the endpoints of every edge with label 1-1 (as such edges are obtained by simplification of vertices with label 1-1); similarly, we can see that $T^{\prime}$ is a tree, as it can be constructed from $T$ by simplification of vertices with label $1$; the final properties can be proved by seeing that $c_{t}(uw)$ corresponds to the maximum congestion in the $u$-$w$-path in $T$, and that $T$ has congestion at most $k$.
As $(S_{t},\ell_{t},c_{t})$ can be constructed from $(S_{t^{\prime}},\ell_{t^{\prime}},c_{t^{\prime}})$ (for introduce and forget nodes) or from $(S_{t_{1}},\ell_{t_{1}},c_{t_{1}})$ and $(S_{t_{2}},\ell_{t_{2}},c_{t_{2}})$ (for join nodes) using the recursion rules, then $D[t,(S_{t},\ell_{t},c_{t})]$ can be obtained from the solutions of the children nodes, and by induction on $t$, $D[t,(S_{t},\ell_{t},c_{t})]$ exists.
Lemma 4.33.
lem:tw-alg:runtime The running time of the algorithm is ${(w+k)}^{{\cal O}(w)}n^{{\cal O}(1)}$.
Proof 4.34.
Let $t\in V(\mathcal{T})$. We start by showing that the number of possible states $(S,\ell,c)$ is ${(w+k)}^{{\cal O}(w)}$. If $(S,\ell)$ is a skeleton, $S$ can have at most $w+1$ vertices with degree at most $2$, since they are in $X_{t}$, and thus it can have at most $w$ vertices with degree at least $3$, for a total of $2w+1$. As a consequence, we can specify $(S,\ell,c)$ by specifying, for each of the at most $2w+1$ vertices in $S$, which vertex is its parent in $S$ ($2w+1$ choices) and its label, and for each edge $e$, its label and value $c(e)\in[k]$. In total, the number of possibilities is upper-bounded by ${(w+k)}^{{\cal O}(w)}$.
The running time of the algorithm is determined by the complexity of recursively calculating each of the ${(w+k)}^{{\cal O}(w)}$ values for the ${\cal O}(nw)$ bags in $(\mathcal{T},\mathcal{X})$. In the case of forget nodes, we can compute $(S,\ell,c)$ from $(S^{\prime},\ell^{\prime},c^{\prime})$ in polynomial time, by updating the congestion values and then applying simplification to the skeleton. Similarly, for introduce nodes, we can compute $(S^{\prime},\ell^{\prime},c^{\prime})$ from $(S,\ell,c)$ in polynomial time, by applying simplification. For join nodes, given states for $t_{1}$, $t_{2}$, we simply have to compute the new labeling and congestion values; we can enumerate the states for $t_{1}$ and $t_{2}$ and compute $(S,\ell,c)$ in time ${(w+k)}^{{\cal O}(w)}$.
Overall, we can compute the solutions for a node $t$ in time ${(w+k)}^{{\cal O}(w)}$, and thus the algorithm runs in time $n\cdot{(w+k)}^{{\cal O}(w)}$.
4.2 FPT Approximation and Clique-width
In this section we leverage the algorithm of Theorem˜4.24 to obtain two further results. On the one hand, we observe that the FPT algorithm of Theorem˜4.24 relies on two parameters (treewidth and the optimal congestion), but it is likely impossible to improve this to an exact algorithm parameterized by treewidth alone (indeed, we saw that Theorem˜1 implies that the problem is hard even for parameters much more restricted than treewidth). We are therefore motivated to consider the question of approximation and present an FPT approximation scheme parameterized by treewidth alone. Our algorithm is based on a combination of the algorithm of Theorem˜4.24 together with a technique introduced in [icalp/Lampis14] (later also used among others in [dmtcs/AngelBEL18, BelmonteLM20, ChuL23, icalp/Lampis21]) which allows us to perform dynamic programming while maintaining approximate values for the congestion. The end result is an FPT approximation scheme, which for any $\varepsilon>0$ is able to return a $(1+\varepsilon)$-approximate solution in time ${(w/\varepsilon)}^{{\cal O}(w)}n^{{\cal O}(1)}$, that is, FPT in $w+\frac{1}{\varepsilon}$.
Having established this, we obtain a second extension of Theorem˜4.24, to the more general parameter clique-width. Here, we rely on a Win/Win argument: suppose we are given an input graph $G$ of clique-width $w$ and are asked if a tree of congestion $k$ can be found; if $G$ contains a large bi-clique (in terms of $k$), then we show that this can be found and we can immediately say No; otherwise, by a well-known result of Gurski and Wanke [wg/GurskiW00], we infer that the graph actually has low treewidth, so we can apply Theorem˜4.24.
Theorem 4.35.
thm:tw-apx There is an algorithm which, for all $\varepsilon>0$, when given as input a graph $G$ of treewidth $w$, returns a spanning tree of congestion at most $(1+\varepsilon)\mathrm{stc}(G)$ in time ${(\frac{w}{\varepsilon})}^{{\cal O}(w)}n^{{\cal O}(1)}$.
Proof 4.36.
We will modify the algorithm of Theorem˜4.24. As before, we assume that we start with a nice tree decomposition of width ${\cal O}(w)$, but now we further apply a result of Bodlaender and Hagerup [siamcomp/BodlaenderH98] which allows us to edit the nice tree decomposition so that it now has height (maximum distance from the root to any leaf) ${\cal O}(w^{2}\log n)$, while the width remains ${\cal O}(w)$. It will allow us to bind the accumulative errors of the dynamic programming algorithm. Let $h$ be the height of the resulting nice tree decomposition.
[…]
In order to modify the algorithm we revisit all the places where a value $c(e)$ is calculated, based on the graph or previously calculated values. This happens in:
[…]
  • Introduce nodes: Operations are unchanged, except we use $\hat{c}$ in the place of $c$.
  • Join nodes: previously we were calculating a triple $(S,\ell,c)$ from two triples $(S,\ell_{1},c_{1})$ and $(S,\ell_{2},c_{2})$, rejecting solutions for which for some edge $e$ we had $c_{1}(e)+c_{2}(e)>k$. Now we reject solutions if $\hat{c}_{1}(e)+\hat{c}_{2}(e)>(1+\varepsilon)k$. If a combination is not rejected, we were previously calculating $c(e)=c_{1}(e)+c_{2}(e)$. Now we set $\hat{c}(e)$ to be the smallest integer power of $(1+\delta)$ that is at least as large as $\hat{c}_{1}(e)+\hat{c}_{2}(e)$, that is, we perform the same addition as before and round-up to the closest integer power of $(1+\delta)$. Other operations are unchanged.
Let us now argue that the algorithm indeed produces a $(1+\varepsilon)$-approximate solution. We initially guess $\mathrm{stc}(G)$ and assume we execute the (modified) algorithm of Theorem˜4.24 for $k=\mathrm{stc}(G)$. We define the height of a bag of the decomposition as the largest distance from the bag to a leaf rooted in its sub-tree. Leaves are at height 0 and the root is at height $h$. Before we proceed, we note that we have set $\delta$ to a value such that ${(1+\delta)}^{h}\leq e^{\delta h}=e^{\frac{\varepsilon}{2}}\leq(1+\varepsilon)$ for small enough $\varepsilon$.
We want to maintain the following invariant:
[…]
Informally, we are claiming here that the approximation algorithm is slightly over-estimating the congestion of every edge: on the one hand, if the approximate algorithm computes some solution, an exact solution with at most the same congestion exists; while on the other hand, if an exact solution exists, the algorithm will compute a congestion upper-bounded by the correct value multiplied by ${(1+\delta)}^{h_{t}}$.
It is now not hard to establish this invariant by induction:
  • For a Forget node $t$ with child $t^{\prime}$: we have $h_{t^{\prime}}=h_{t}-1$. By induction, for an exact triple $(S^{\prime},\ell^{\prime},c^{\prime})$ we have calculated an approximate triple $(S^{\prime},\ell^{\prime},\hat{c}^{\prime})$ with $\hat{c}^{\prime}(e)\leq{(1+\delta)}^{h_{t}-1}c^{\prime}(e)$. For each edge $e$ the exact triple $(S,\ell,c)$ would contain a value $c(e)=c^{\prime}(e)+r_{e}$, while we set $\hat{c}(e)\leq(1+\delta)(\hat{c}^{\prime}(e)+r_{e})\leq{(1+\delta)}^{h_{t}}c(e)$, as desired. For the other direction, we observe that we set $\hat{c}(e)\geq\hat{c}^{\prime}(e)+r_{e}$, so using that by induction $\hat{c^{\prime}}(e)\geq c^{\prime}(e)$ we get the second statement of the invariant. We observe that discarding a solution is correct, because for all valid triples calculated by the exact algorithm, for each value $c(e)\leq k$, an approximate triple will have value at most $\hat{c}(e)\leq{(1+\delta)}^{h}c(e)\leq(1+\varepsilon)k$, hence solutions with $\hat{c}(e)>(1+\varepsilon)k$ are safe to discard.
  • For Introduce nodes, congestion values are not modified, but the height increases, so the invariant remains true.
  • For Join nodes, where a node $t$ has two children $t_{1},t_{2}$, we have $h_{t_{1}},h_{t_{2}}\leq h_{t}-1$. We use similar reasoning as in the Forget case, in particular, we compute $\hat{c}(e)\leq(1+\delta)(\hat{c}_{1}(e)+\hat{c}_{2}(e))$, but by inductive hypothesis $\hat{c}_{1}(e)\leq{(1+\delta)}^{h_{t}-1}c_{1}(e)$ (and similarly for $\hat{c}_{2}$), so $\hat{c}(e)$ over-estimates the correct value by at most a factor ${(1+\delta)}^{h_{t}}$. Similarly, $\hat{c}(e)\geq\hat{c}_{1}(e)+\hat{c}_{2}(e)$, so if $\hat{c}_{1}(e)\geq c_{1}(e)$ (and similarly for $\hat{c}_{2}$) we get the second statement. Again, discarding a solution with congestion strictly greater than $(1+\varepsilon)k$ is justified as in the Forget case.
It is not hard to see that the invariant guarantees that if a solution exists the approximation algorithm will find it and that if the approximation algorithm finds a solution of congestion estimated at $\hat{k}\leq(1+\varepsilon)k$, a solution of this congestion can be constructed. Let us then bound the running time, which is dominated by the size of the DP tables. As before, there are $w^{{\cal O}(w)}$ trees $S$, $2^{{\cal O}(w)}$ functions $\ell$, while the number of possible values stored for $\hat{c}(e)$ is at most ${\cal O}(\log_{(1+\delta)}k)={\cal O}(\frac{\log k}{\log(1+\delta)})={\cal O}(\frac{\log n}{\delta})={\cal O}(\frac{w^{2}\log^{2}n}{\varepsilon})$, where we used that $h={\cal O}(w^{2}\log n)$ and approximated $\log(1+\delta)$ by $\delta$. The total number of possible DP entries is then ${(\frac{w\log n}{\varepsilon})}^{{\cal O}(w)}$. We now use a standard Win/Win argument: if $w\leq\sqrt{\log n}$, then the running time bound we calculated is $n^{o(1)}$ and the whole algorithm runs in polynomial time; otherwise, $\log n<w^{2}$, so the running time of the algorithm is ${(\frac{w}{\varepsilon})}^{{\cal O}(w)}n^{{\cal O}(1)}$.
Theorem 4.37.
thm:cw There is an algorithm which, when given as input a graph $G$, an integer $k$, and a clique-width expression for $G$ with $w$ labels, correctly decides if $\mathrm{stc}(G)\leq k$ in time ${(wk)}^{{\cal O}(wk)}n^{{\cal O}(1)}$.
Proof 4.38.
We use the following claim:
Claim 9.
If, for some $t$, a graph $G$ has $K_{t,t}$ as a subgraph, then $\mathrm{stc}(G)\geq t$.
{claimproof}
Theorem 1(a) in [dm/Ostrovskii04] states that $\mathrm{stc}(G)\geq m_{G}$”, where $m_{G}$ is the maximum number of edge-disjoint paths between two vertices in $G$. We can see that $m_{K_{t,t}}=t$. Hence, if a graph $G$ contains $K_{t,t}$ as a subgraph, $\mathrm{stc}(G)\geq t$.
The algorithm now relies on a result of [wg/GurskiW00] which states that if a graph has clique-width $w$ and does not contain $K_{t,t}$ as a subgraph, then $G$ has treewidth at most $3tw$. Given $G$, $k$, and a clique-width expression with $w$ labels, our algorithm first checks if $G$ has treewidth at most $3(k+1)w$ by running the 2-approximation of [Korhonen21]. If this algorithm does not return a tree decomposition of width at most $6(k+1)w+1$, we can immediately conclude that the treewidth of $G$ is more than $3(k+1)w$. However, if $G$ did not contain any bicliques $K_{k+1,k+1}$, this could not happen, so $G$ must contain such a biclique, which by the claim allows us to immediately return No. We are therefore left with the case where we have a tree decomposition of width ${\cal O}(kw)$, on which we execute the algorithm of Theorem˜4.24.
5 FPT Algorithms
Given the hardness results of Section˜3, we are motivated to search for structural parameters that render Spanning Tree Congestion fixed-parameter tractable. In this section we present various such results, starting with the parameter distance to clique in Section˜5.1, then vertex integrity in Section˜5.2, and finally feedback edge number in Section˜5.3.
5.1 Distance to Clique
Corollary˜3.3 implies that Spanning Tree Congestion remains W[1]-hard even on very structured dense instances. In this subsection we search for parameters that render the problem tractable on dense instances, and present an FPT algorithm when parameterized by the distance to clique of the input graph, arguably one of the most restrictive such parameters. Interestingly, the running time of our algorithm is dictated by the “easy” case, where the clique modulator is large with respect to the size of the graph. We remark that a modulator to clique of a graph $G$ is a vertex cover in the complement $\bar{G}$ of $G$, and thus the minimum modulator can be computed by employing any FPT algorithm for Vertex Cover (e.g. [tcs/ChenKX10, stacs/HarrisN24]) on $\bar{G}$.
Theorem 5.39.
thm:distance_to_clique Given $G=(V,E)$ and $S\subseteq V$ with GSG-S being a clique, there is an algorithm that returns a spanning tree of $G$ of congestion $\mathrm{stc}(G)$ in time $2^{{\cal O}(k^{3})}n^{{\cal O}(1)}$, where $n=|V|$ and $k=|S|$.
Proof 5.40.
Let $C=V\setminus S$ denote the vertex set of the clique GSG-S, with $c=|C|$. First consider the case where $c\leq 2k^{3}+4k={\cal O}(k^{3})$. Then, it holds that $n=c+k={\cal O}(k^{3})$, and one can find a spanning tree of $G$ of congestion $\mathrm{stc}(G)$ in time $2^{{\cal O}(k^{3})}n^{{\cal O}(1)}$ by using the $2^{n}n^{{\cal O}(1)}$ algorithm of Okamoto, Otachi, Uehara, and Uno [jgaa/OkamotoOUU11].
For the remainder of the proof assume that $c>2k^{3}+4k$. Given a spanning tree $T$ of $G$ and an edge $e=\{u_{1},u_{2}\}\in E(T)$, let $V_{u_{1}}\uplus V_{u_{2}}=V$ denote the subsets of vertices of $V$ that belong to the connected component of TeT-e containing vertices $u_{1}$ and $u_{2}$ respectively; notice that $V_{u_{1}},V_{u_{2}}$ define a partition on $V$. The crux of our algorithm is the following claim, which allows us to upper-bound the value of $\mathrm{stc}(G)$.
Claim 10.
There exists a spanning tree $T$ of $G$ with $\mathrm{cng}_{G}(T)<2c-\frac{c}{k}+2k^{2}$.
{claimproof}
Let $S^{\prime}=\{s\in S\mid|N_{G}(s)\cap C|>c-c/k\}\subseteq S$ denote the set of vertices of $S$ with more than cc/kc-c/k neighbors in $C$. It holds that there exists $r\in C$ such that $S^{\prime}\subseteq N_{G}(r)$. This follows by the pigeonhole principle, as the number of vertices in $C$ having at least one non-neighbor in $S^{\prime}$ is at most $\sum_{s\in S^{\prime}}|C\setminus N_{G}(s)|<\sum_{s\in S^{\prime}}c/k\leq c$. Let $S_{0}=N_{G}(r)\cap S$ denote the neighbors of $r$ in $S$, where $S_{0}\supseteq S^{\prime}$.
We define $G^{\prime}=(V,E^{\prime})$ to be the subgraph of $G$ on edge set $E^{\prime}=E(C\cup S_{0},S\setminus S_{0})$, and let $M\subseteq E^{\prime}$ denote a maximum matching in $G^{\prime}$, with $S_{M}\subseteq S\setminus S_{0}$ denoting the set of vertices of $S\setminus S_{0}$ that take part in $M$. Notice that for any $s\in S_{M}$, it holds that $|N_{G}(s)\cap C|\leq c-\frac{c}{k}$, as $S_{M}\cap S^{\prime}=\varnothing$.
To construct the spanning tree $T$, we first include all edges incident to $r$ in $E(T)$ and then add all edges present in $M$. Let $Z=S\setminus(S_{0}\cup S_{M})$. To conclude the construction of $T$, arbitrarily include some edges incident to the vertices of $Z$ in order to obtain a spanning tree.
It remains to show that $\mathrm{cng}_{G,T}(e)<2c-\frac{c}{k}+2k^{2}$ for all $e=\{u_{1},u_{2}\}\in E(T)$. We consider different cases, depending on whether $r\in e$ or not. Before doing so, notice that for any such edge, it holds that
$\mathrm{cng}_{G,T}(e)\leq\min\left\{\sum_{u\in V_{u_{1}}}\deg_{G}(u),\sum_{u\in V_{u_{2}}}\deg_{G}(u)\right\}.$
Additionally, we remark that $\forall z\in Z$, $|N_{G}(z)\cap C|\leq k$, which implies that $\deg_{G}(z)<2k$. To see this, notice that since $M$ is a maximum matching, no vertex in $Z$ is adjacent with a vertex of $C\cup S_{0}$ that does not take part in $M$. Thus, $|N_{G}(z)\cap C|\leq|M|\leq|S|\leq k$.
First consider the case where $r\notin e$. Then, it holds that either $C\subseteq V_{u_{1}}$ or $C\subseteq V_{u_{2}}$. Assume without loss of generality that $C\subseteq V_{u_{1}}$. In that case, notice that $V_{u_{2}}\cap S_{0}=\varnothing$, while $|V_{u_{2}}\cap S_{M}|\leq 1$ and any other vertex of $V_{u_{2}}$ belongs to $Z$. Consequently, it follows that
$\mathrm{cng}_{G,T}(e)\leq\sum_{u\in V_{u_{2}}}\deg_{G}(u)\leq(c-\frac{c}{k}+k)+(k-1)\cdot 2k<c-\frac{c}{k}+2k^{2},$
with the first term being due to the at most one vertex of $V_{u_{2}}\cap S_{M}$ and the second due to the vertices of $V_{u_{2}}\cap Z$.
Alternatively, $e=\{r,w\}$, where $w\in S_{0}\cup(C\setminus\{r\})$. Notice that in this case, it holds that $|V_{w}\cap S_{M}|\leq 1$, while any other vertex of $V_{w}$ belongs to $Z$. Consequently,
$\mathrm{cng}_{G,T}(e)\leq\sum_{u\in V_{w}}\deg_{G}(u)<(c+k)+(c-\frac{c}{k}+k)+(k-2)\cdot 2k<2c-\frac{c}{k}+2k^{2},$
with the first term being due to $\deg_{G}(w)$, the second due to the at most one vertex of $V_{w}\cap S_{M}$, and the third due to the vertices of $V_{w}\cap Z$.
The previous claim now allows us to determine the structure of a spanning tree $T_{\operatorname{OPT}}$ of $G$ of minimum congestion.
Claim 11.
For any $e=\{u_{1},u_{2}\}\in E(T_{\operatorname{OPT}})$ it holds that either $|V_{u_{1}}\cap C|\leq 1$ or $|V_{u_{2}}\cap C|\leq 1$.
{claimproof}
Let $e=\{u_{1},u_{2}\}\in E(T_{\operatorname{OPT}})$. By Claim˜10 and the fact that $c>2k^{3}+4k$ it follows that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)<2c-\frac{c}{k}+2k^{2}<2c-2k^{2}-4+2k^{2}=2(c-2)$. Assume that $|V_{u_{1}}\cap C|=x$ and $|V_{u_{2}}\cap C|=c-x$, where $x\in[2,c-2]$. In that case, it holds that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)\geq x\cdot(c-x)$ due to the vertices of $C$. Since $x\cdot(c-x)\geq 2\cdot(c-2)$ for all $x\in[2,c-2]$, it follows that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)\geq 2(c-2)$, a contradiction.
Consequently, due to Claim˜11 it follows that there exists a vertex $r\in V$ such that in $T_{\operatorname{OPT}}-r$, every connected component contains at most one vertex of $C$. Notice that the vertices of $S$ are present in at most $k$ such connected components. Therefore, in order to compute $T_{\operatorname{OPT}}$, we initially guess vertex $r$, and due to the indistinguishability of twin vertices, this amounts to a total of $2^{k}+k$ choices. Next, we guess the set $Q\subseteq C$, which contains the vertices of $C$ present in connected components of $T_{\operatorname{OPT}}-r$ that contain vertices of $S$. Since $|Q|\leq k$ and vertices belonging to the same twin classes are indistinguishable, there are a total of $\sum_{i=0}^{k}{(2^{k})}^{i}\leq k\cdot 2^{k^{2}}$ choices regarding the set $Q$. For the rest of the vertices $c\in C\setminus Q$, we add the edge $\{r,c\}$ in $E(T_{\operatorname{OPT}})$ (if such an edge does not exist, our guess is wrong and we discard it). Finally, since $T_{\operatorname{OPT}}$ is a tree, it suffices to guess the parent for each vertex in $S\cup Q$, among the set $S\cup Q\cup\{r\}$, which gives a total of at most ${(2k+1)}^{2k}$ possibilities. Given one such spanning tree, one can in polynomial time compute its congestion. The final running time is $2^{{\cal O}(k^{2})}n^{{\cal O}(1)}$.
5.2 Vertex integrity
Here we prove that the parameterization by vertex integrity renders Spanning Tree Congestion FPT.
Theorem 5.41.
thm:vi Spanning Tree Congestion is fixed-parameter tractable parameterized by vertex integrity.
Proof 5.42.
We say that two spanning trees $T$ and $T^{\prime}$ of $G=(V,E)$ are equivalent if there is an automorphism $\eta$ of $G$ that is an isomorphism from $T$ to $T^{\prime}$ as well. Observe that the detour in $T$ for $\{u,v\}\in E(G)$ contains $\{x,y\}\in E(T)$ if and only if the detour in $T^{\prime}$ for $\{\eta(u),\eta(v)\}\in E(G)$ contains $\{\eta(x),\eta(y)\}\in E(T^{\prime})$. This implies that equivalent spanning trees have the same congestion.
{observation}
If $T$ and $T^{\prime}$ are equivalent spanning trees of $G$, then $\mathrm{cng}_{G}(T)=\mathrm{cng}_{G}(T^{\prime})$.
[…]
*Signatures of spanning trees. Let $k=\mathrm{vi}(G)$ and $S\subseteq V(G)$ be a vertex set such that $|S|+\max_{C\in\mathtt{cc}(G-S)}|V(C)|\leq k$. We say that $C,C^{\prime}\in\mathtt{cc}(G-S)$ have the same type if there is an isomorphism $\eta$ from $C$ to $C^{\prime}$ such that $N(v)\cap S=N(\eta(v))\cap S$ for every $v\in V(C)$. We call such an isomorphism a type isomorphism. Note that the relation of having the same type is an equivalence relation on $\mathtt{cc}(G-S)$. Let $\mathcal{C}_{1},\dots,\mathcal{C}_{p}$ be the equivalence classes defined by this relation. That is, each $\mathcal{C}_{i}$ is a maximal subset of $\mathtt{cc}(G-S)$ with the same type. For the number $p$ of different types it holds that $p\leq 2^{k^{2}}$, as there are at most $2^{k^{2}}$ non-isomorphic labeled graphs with at most $k$ vertices.
Let $C,C^{\prime}\in\mathcal{C}_{i}$ for some $i\in[p]$. Let $F$ and $F^{\prime}$ be spanning forests of $C_{S}\coloneqq G[V(C)\cup S]$ and $C^{\prime}_{S}\coloneqq G[V(C^{\prime})\cup S]$, respectively. We say that $F$ and $F^{\prime}$ have the same forest type if there is an isomorphism $\eta$ from $C_{S}$ to $C^{\prime}_{S}$ satisfying the following conditions:
  • $\eta(v)=v$ for every $v\in S$;
  • $\eta$ is an isomorphism from $F$ to $F^{\prime}$;
  • $\eta$ restricted to $V(C)$ is a type isomorphism from $C$ to $C^{\prime}$.
For $i\in[p]$, let $q_{i}$ be the number of different forest types for $\mathcal{C}_{i}$. This number can be upper-bounded by the number of possible ways of selecting at most k1k-1 edges from at most $\binom{k}{2}$ edges, and thus $q_{i}\leq k^{2k}$ holds. We fix an ordering on the different forest types of such spanning forests in an arbitrary way. This allows us to specify a forest type as “the $j$-th forest type of $\mathcal{C}_{i}$” for example.
Let $T$ be a spanning tree of $G$. A component $C$ in $\mathtt{cc}(G-S)$ has type $(i,j)$ in $T$ if $C$ belongs to $\mathcal{C}_{i}$ and $T[V(C)\cup S]$ has the $j$-th forest type of $\mathcal{C}_{i}$. The signature of $T$ is the pair of $T[S]$ and the mapping $\sigma_{T}$ such that $\sigma_{T}(i,j)$ is the number of type $(i,j)$ components in $T$. We show that having the same signature implies the equivalence of two spanning trees.
Claim 12.
If spanning trees $T$ and $T^{\prime}$ of $G$ have the same signature, then $\mathrm{cng}_{G}(T)=\mathrm{cng}_{G}(T^{\prime})$.
{claimproof}
Due to ˜5.42, it suffices to show that $T$ and $T^{\prime}$ are equivalent. We show that there is an automorphism $\eta$ of $G$ that is an isomorphism from $T$ to $T^{\prime}$. Let $(F,\sigma)$ be the common signature of $T$ and $T^{\prime}$. For all $v\in S$, set $\eta(v)=v$. Let $i\in[p]$ and $j\in[q_{i}]$. We order the components of type $(i,j)$ in $T$ in an arbitrary way as $C_{1},\dots,C_{\sigma(i,j)}$. Similarly, we order the ones in $T^{\prime}$ as $C^{\prime}_{1},\dots,C^{\prime}_{\sigma(i,j)}$. For each $h\in[\sigma(i,j)]$, we extend $\eta$ by mapping $C_{h}$ to $C^{\prime}_{h}$ with the type isomorphism certifying that they are of type $(i,j)$. The definition of being type $(i,j)$ ensures that the resulting mapping $\eta$ is an automorphism of $G$ and an isomorphism from $T$ to $T^{\prime}$.
[…]
Leaf components and non-leaf components in a spanning tree. Let $T$ be a spanning tree of $G$. We call $C\in\mathtt{cc}(G-S)$ a non-leaf component in $T$ if there exists $C^{\prime}\in\mathtt{cc}(T[V(C)])$ such that $|N_{T}(V(C^{\prime}))\cap S|\geq 2$; otherwise, we call $C$ a leaf component in $T$. Intuitively, a non-leaf component connects some vertices in $S$, while a leaf component does not. Observe that from the type $(i,j)$ in $T$ of a component, we can easily check if the component is a leaf or a non-leaf.
We can see that the number of non-leaf components is small as $T$ contains no cycle.
{observation}
There are at most |S|1|S|-1 non-leaf components in $T$.
Let $H$ be the subgraph of $G$ induced by $S$ and the vertices in non-leaf components. Observe that adding leaf components to $H$ never decreases the number of connected components in $T[V(H)]$. Hence, the following holds.
{observation}
$T[V(H)]$ is a spanning tree of $H$. Furthermore, for each leaf component $C$ in $T$, $T[V(H)\cup V(C)]$ is a spanning tree of $G[V(H)\cup V(C)]$.
By ˜5.42, an edge in $H$ is detoured in $T[V(H)]$ and an edge in a leaf component $C$ is detoured in $T[V(H)\cup V(C)]$. This implies the following locality of leaf components.
{observation}
Let $C$ be a leaf component in $T$ and $e$ be an edge incident to a vertex in $C$. If $e\in E(T)$, then every edge using $e$ in its detour in $T$ is incident to a vertex in $C$. If $e\notin E(T)$, then the detour for $e$ in $T$ contains an edge $f$ if and only if the detour for $e$ in $T[V(H)\cup V(C)]$ contains $f$.
By ˜5.42, we can see that in any spanning tree $T$, an edge incident to a vertex of a leaf component $C$ has congestion at most $|E(S\cup V(C))|<k^{2}$. Therefore, if we know that $\mathrm{stc}(G)\geq k^{2}$, then we only have to bound the congestion of the edges in $T[V(H)]$. This will be helpful for simplifying the description of the algorithm below. In fact, we can assume that $\mathrm{stc}(G)\geq k^{2}$ since we can check whether $\mathrm{stc}(G)=s$ for every $s<k^{2}$ by using a fixed-parameter algorithm parameterized by $\mathrm{tw}(G)+\mathrm{stc}(G)$ (e.g., the one of Theorem˜4.24).
[…]
The algorithm. We are ready to describe the algorithm, which can be outlined as follows.
  1. 1.
    Guess the non-leaf components in an optimal spanning tree $T$.
    • By ˜5.42, there are at most |S|1|S|-1 non-leaf components in $T$.
  2. 2.
    Let $H$ be the subgraph of $G$ induced by $S$ and the vertices in non-leaf components. Guess $T_{H}\coloneqq T[V(H)]$.
    • By ˜5.42, $T_{H}$ is a spanning tree of $H$.
  3. 3.
    Find the rest of the tree $T$.
    • By Claim˜12, it suffices to find an optimal signature of $T$ consistent with the guesses.
    • Since all remaining components are leaf components, ˜5.42 allows us to use integer linear programming (ILP) for this task.
As discussed above, we assume that $\mathrm{stc}(G)\geq k^{2}$, and thus we only need to upper-bound the congestion of the edges in $T_{H}$.
For the non-leaf components, we have at most $p^{k}\leq 2^{k^{3}}$ possibilities for selecting at most k1k-1 components from $p$ classes. For $i\in[p]$, let $\mathcal{C}^{\prime}_{i}$ be the set of components obtained from $\mathcal{C}_{i}$ by removing the selected components (if no component is selected from $\mathcal{C}_{i}$, then $\mathcal{C}^{\prime}_{i}=\mathcal{C}_{i}$). Given the non-leaf components, there are at most $k^{2k^{2}}$ possibilities for $T_{H}$ as $H$ has at most $k^{2}$ vertices and $T_{H}$ is a spanning tree of $H$. Thus, in total, there are at most $2^{k^{3}}\cdot k^{2k^{2}}$ candidates for $T_{H}$. Therefore, by paying the multiplicative factor of $2^{k^{3}}\cdot k^{2k^{2}}$ in the running time, we can assume that $T_{H}$ is correctly guessed.
Now we find the rest of the tree $T$ using ILP (the third step in the outline). For each $i\in[p]$ and each $j\in[q_{i}]$, let $x_{i,j}$ be the non-negative integer variable representing the value $\sigma_{T}(i,j)$; that is, it determines the number of components of type $(i,j)$ in $T$. If components of type $(i,j)$ are non-leaf in $T$, then we set $x_{i,j}=0$. For all $i\in[p]$, we set $\sum_{j\in[q_{i}]}x_{i,j}=|\mathcal{C}^{\prime}_{i}|$ to make the number of used components in $\mathcal{C}^{\prime}_{i}$ correct.
For each edge $e$ in $T_{H}$, take the positive integer variable $y_{e}$ that represents the congestion of $e$ in $T$. Let $a_{e}$ be the number of edges in $H$ that contribute to the congestion of $e$ in $T$. Let $b_{(i,j),e}$ be the number edges in a component of type $(i,j)$ in $T$ that contribute to the congestion of $e$ in $T$. By ˜5.42, we can compute these constants in advance. Now we can represent the congestion of $e$ in $T$ by setting $y_{e}=a_{e}+\sum_{i\in[p],\;j\in[q_{i}]}b_{(i,j),e}\cdot x_{i,j}$.
Finally, we set the objective of the formula to minimizing the maximum of $y_{e}$. That is, we introduce a variable $z$ and put the constraint $y_{e}\leq z$ for all $e\in E(H)$, and then set $z$ as the objective function to be minimized.
The number of variables in the formula is at most $p\cdot\max_{i\in[p]}q_{i}+|E(H)|+1$, which is ${\cal O}(2^{k^{2}}k^{2{k}})$. It is known that ILP parameterized by the number of variables is fixed-parameter tractable [mor/Lenstra83], and thus we can conclude that our algorithmic task here is fixed-parameter tractable parameterized by $k$, the vertex integrity of $G$. This implies the theorem. (Note that using the recent result of Reis and Rothvoss [focs/ReisR23], ILP with $d$ variables can be solved in randomized ${(\log(2d))}^{{\cal O}(d)}$ time. Hence, this part can be done in randomized $k^{{\cal O}(2^{k^{2}}k^{2{k}})}$ time.)
5.3 Feedback Edge Number
Notice that the algorithm of Theorem˜4.24 already implies that Spanning Tree Congestion is FPT parameterized by the feedback edge number $\mathrm{fes}$ of the input graph: any instance with $k\geq\mathrm{fes}+1$ is trivially yes, thus one can easily obtain an $\mathrm{fes}^{{\cal O}(\mathrm{fes})}n^{{\cal O}(1)}$ algorithm. A natural question is whether the slightly super-exponential parametric dependence can be overcome, and we answer this affirmatively by providing such an algorithm. In fact, more strongly, we present a kernelization algorithm that results in a graph with only ${\cal O}(\mathrm{fes})$ vertices and edges, thus allowing us to exhaustively guess the spanning tree of minimum congestion. To do so, we notice that we can safely delete vertices of degree $1$, resulting in a graph with only ${\cal O}(\mathrm{fes})$ vertices of degree larger than $2$. Next, we can contract most of the remaining edges, thereby leaving only ${\cal O}(\mathrm{fes})$ of them, thus allowing us to guess exhaustively which edges belong to an optimal solution.
Theorem 5.43.
Spanning Tree Congestion admits a kernel with ${\cal O}(\mathrm{fes})$ vertices and edges, where $\mathrm{fes}$ denotes the feedback edge number of the input graph.
Proof 5.44.
Let $G_{0}$ denote the input connected graph. We start by exhaustively deleting vertices of degree $1$ in $G_{0}$; it is easy to see that this is safe, as any such vertex is connected to any spanning tree via the single edge it is incident with, which is of congestion $1$. Let $G$ denote the resulting connected graph, where $\mathrm{fes}$ denotes its feedback edge number; notice that the deletion of vertices cannot increase the feedback edge number of a graph. Let $V^{=2}(G)$ and $V^{\geq 3}(G)$ denote the sets of vertices of $G$ of degree exactly $2$ and at least $3$ respectively; notice that they induce a partition on $V(G)$, since all vertices of $G$ are of degree at least $2$. Bentert, Dittmann, Kellerhals, Nichterlein, and Niedermeier [jgaa/BentertDKNN20, Lemma 2] have proved that $|V^{\geq 3}(G)|<2\mathrm{fes}$ and $\sum_{v\in V^{\geq 3}(G)}\deg_{G}(v)=2(|V^{\geq 3}(G)|+\mathrm{fes}-1)<6\mathrm{fes}$. [^7] We mention in passing a slight mistake in their proof, where they claim that $\sum_{v\in V^{\geq 3}(G)}\deg_{G}(v)=3|V^{\geq 3}(G)|$ instead. We next define the following reduction rule.
[…]
*Rule $(\diamond)$. Let $G$ be a graph with (not necessarily distinct) vertices $u,v\in V^{\geq 3}(G)$ such that there is a uuvv path $P$ in $G$ whose internal vertices all belong to $V^{=2}(G)$. Let $L_{P}\geq 1$ be equal to the number of internal vertices of $P$. Then, {romanenumerate}
if $u=v$ and $L_{P}>2$, contract edges in $P$ such that only $2$ internal vertices are left,
if $u\neq v$, $\{u,v\}\in E(G)$, and $L_{P}>1$, contract edges in $P$ such that only $1$ internal vertex is left,
if $u\neq v$ and $\{u,v\}\notin E(G)$, delete all internal vertices of $P$ and add the edge $\{u,v\}$.
Notice that Rule $(\diamond)$ can be applied at most $n$ times, since each of its applications reduces the number of vertices of the graph. Let $G^{\prime}$ denote the connected graph obtained after exhaustively doing so. Notice that $G$ can be obtained from $G^{\prime}$ by only subdividing edges; edge subdivision does not change the spanning tree congestion of unweighted graphs [algorithmica/BodlaenderFGOL12, Lemma 7.10], thus it holds that $\mathrm{stc}(G)=\mathrm{stc}(G^{\prime})$. Furthermore, it is easy to see that $V^{\geq 3}(G^{\prime})=V^{\geq 3}(G)$, and in fact any such vertex has the same degree in both $G$ and $G^{\prime}$. Consequently, it follows that $\sum_{v\in V^{\geq 3}(G^{\prime})}\deg_{G^{\prime}}(v)=\sum_{v\in V^{\geq 3}(G)}\deg_{G}(v)<6\mathrm{fes}$.
Let $E_{1}$ and $E_{2}$ define a partition on the edges of $G^{\prime}$, where $E_{1}=\{e\in E(G^{\prime})\mid e\cap V^{\geq 3}(G^{\prime})\neq\varnothing\}$ denotes the set of edges in $E(G^{\prime})$ with at least one endpoint belonging to $V^{\geq 3}(G^{\prime})$, and $E_{2}$ the rest, whose endpoints both belong to $V^{=2}(G^{\prime})$. Due to the previous discussion, it holds that $|E_{1}|<6\mathrm{fes}$. As for $E_{2}$, since Rule $(\diamond)$ has been applied exhaustively, any such edge can only be due to Case (5.44) of Rule $(\diamond)$, thus $|E_{2}|\leq|E_{1}|/2$ holds. In total, it follows that $|E(G^{\prime})|<9\mathrm{fes}$. Furthermore, since $G^{\prime}$ is connected, it follows that $|V(G^{\prime})|\leq|E(G^{\prime})|+1$.
6 Conclusion
In this paper we have presented a number of new results on the parameterized complexity of Spanning Tree Congestion, painting an almost-complete picture regarding its tractability under the most standard parameterizations. As a direction for future work, it would be interesting to consider the problem’s tractability parameterized by the neighborhood diversity, the treewidth plus the maximum degree, as well as whether an FPT-AS parameterized by clique-width exists. Furthermore, although the problem is FPT by vertex cover due to LABEL:thm:vi, the algorithm is based on an ILP, and a simpler (and faster) combinatorial algorithm might be possible under this parameterization; along these lines, it would be interesting to determine whether the problem admits a polynomial kernel in this case. We additionally mention that it is unknown whether the problem remains NP-hard on cographs or when $k=4$.